### v1

# 37N35, 65F55, 93A15, 93B15, 93C57

Abstract
Structured reduced-order modeling is a central component in the computer-aided design of control systems in which cheap-to-evaluate low-dimensional models with physically meaningful internal structures are computed. In this work, we develop a new approach for the structured data-driven surrogate modeling of linear dynamical systems described by second-order time derivatives via balanced truncation model-order reduction. The proposed method is a data-driven reformulation of position-velocity balanced truncation for second-order systems and generalizes the quadrature-based balanced truncation for unstructured first-order systems to the second-order case. The computed surrogates encode a generalized proportional damping structure, and the damping coefficients are inferred solely from data by minimizing a least-squares error over the coefficients. Several numerical examples demonstrate the effectiveness of the proposed method.
keywords:
data-driven modeling, reduced-order modeling, second-order systems, mechanical systems, balanced truncation, transfer function data
[…]
1 Introduction
The mathematical modeling of, e.g., mechanical or electrical structures [45, Ch. 1], typically results in linear dynamical systems described by second-order differential equations of the form
$\mathcal{G}:\left\{\begin{aligned} \boldsymbol{M}\ddot{\boldsymbol{x}}(t)+\boldsymbol{D}\dot{\boldsymbol{x}}(t)+\boldsymbol{K}\boldsymbol{x}(t)&=\boldsymbol{B}_{\operatorname{u}}\boldsymbol{u}(t),\quad\boldsymbol{x}(0)=\dot{\boldsymbol{x}}(0)=\boldsymbol{0},\\ \boldsymbol{y}(t)&=\boldsymbol{C}_{\operatorname{p}}\boldsymbol{x}(t)+\boldsymbol{C}_{\operatorname{v}}\dot{\boldsymbol{x}}(t),\end{aligned}\right.$ (1)
where $\boldsymbol{M},\boldsymbol{D},\boldsymbol{K}\in\mathbb{R}^{n\times n}$ model the internal system dynamics, $\boldsymbol{B}_{\operatorname{u}}\in\mathbb{R}^{n\times m}$ describes the external inputs and $\boldsymbol{C}_{\operatorname{p}},\boldsymbol{C}_{\operatorname{v}}\in\mathbb{R}^{p\times n}$ model the quantities of interest. The three matrices $\boldsymbol{M}$, $\boldsymbol{D}$ and $\boldsymbol{K}$ are typically referred to as mass, damping and stiffness terms due to their physical interpretation. The internal behavior of the system eq. 1 is given by the states $\boldsymbol{x}\colon[0,\infty)\to\mathbb{R}^{n}$, which can be externally influenced by the inputs $\boldsymbol{u}\colon[0,\infty)\to\mathbb{R}^{m}$ and the quantities of interest are observed as outputs $\boldsymbol{y}\colon[0,\infty)\to\mathbb{R}^{p}$. An equivalent description of the input-to-output response of eq. 1 is given in the frequency domain via the corresponding transfer function
$\boldsymbol{G}(s)=\left(s\boldsymbol{C}_{\operatorname{v}}+\boldsymbol{C}_{\operatorname{p}}\right)\left(s^{2}\boldsymbol{M}+s\boldsymbol{D}+\boldsymbol{K}\right)^{-1}\boldsymbol{B}_{\operatorname{u}},$ (2)
with the complex variable $s\in\mathbb{C}$. Systems with such second-order internal structure eq. 1 arise in a variety of applications, ranging from the vibrational analysis of mechanical and acoustical structures [45, 4, 48] to the modeling of electro-mechanical systems [13, 14].
Due to a generally high demand for modeling accuracy, mathematical models such as eq. 1 are typically described by a large number of differential equations, $n\in\mathcal{O}(n^{6})$, making their use in downstream computational tasks such as numerical simulation or optimization very challenging. Simultaneously, in complex applications, explicit state-space models of the form eq. 1 may be either difficult to obtain or wholly unavailable. Instead, the underlying system can only be accessed in the form of data. This motivates data-driven reduced-order modeling, which is the construction of low-dimensional surrogate models solely from system data. In this work, we consider the identification of structured second-order systems eq. 1 directly from input-to-output data. In other words, our goal is to learn a reduced-order model of the form
$\widetilde{\mathcal{G}}:\left\{\begin{aligned} {\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}\ddot{{\widetilde{\boldsymbol{x}}\mkern 2.0mu\mkern-2.0mu}{}}(t)+{\widetilde{\boldsymbol{D}\mkern 2.0mu}\mkern-2.0mu}{}\dot{{\widetilde{\boldsymbol{x}}\mkern 2.0mu\mkern-2.0mu}{}}(t)+{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}{\widetilde{\boldsymbol{x}}\mkern 2.0mu\mkern-2.0mu}{}(t)&={\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\boldsymbol{u}(t),\quad{\widetilde{\boldsymbol{x}}\mkern 2.0mu\mkern-2.0mu}{}(0)=\dot{{\widetilde{\boldsymbol{x}}\mkern 2.0mu\mkern-2.0mu}{}}(0)=\boldsymbol{0},\\ \widetilde{\boldsymbol{y}}(t)&={\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}}{\widetilde{\boldsymbol{x}}\mkern 2.0mu\mkern-2.0mu}{}(t)+{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}\dot{{\widetilde{\boldsymbol{x}}\mkern 2.0mu\mkern-2.0mu}{}}(t),\end{aligned}\right.$ (3)
from evaluations of the transfer function eq. 2, with the reduced-order quantities ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{D}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\in\mathbb{R}^{r\times r}$, ${\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\in\mathbb{R}^{r\times m}$, ${\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}},{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}\in\mathbb{R}^{p\times r}$, ${\widetilde{\boldsymbol{x}}\mkern 2.0mu\mkern-2.0mu}{}\colon[0,\infty)\to\mathbb{R}^{r}$, and $\widetilde{\boldsymbol{y}}\colon[0,\infty)\to\mathbb{R}^{p}$, and a much smaller number of differential equations $r\ll n$.
Second-order systems eq. 1 can always be recast in unstructured (first-order) form by introducing auxiliary variables. Thus, it is in principle always possible to apply classical data-driven modeling techniques to realize unstructured models of the underlying dynamics in eq. 1. However, while it is always possible to write second-order systems in first-order form, the converse is in general not true [26, Sec. III]. Thus, when using unstructured data-driven modeling methods, the underlying internal system structure is typically not retrievable. This is usually an undesired outcome because the internal system structure is what enables the physical reinterpretation of the learned terms, allows for the use of established computational tools, and preserves structure-inherent properties. Further, twice the number of degrees of freedom is required to realize a first-order system that recovers the input-to-output response of the second-order dynamics, and so structured surrogates typically produce more accurate approximations compared to unstructured ones having the same model order. We refer the reader to [45, 37] for comparisons of methods for structured and non-structured surrogate modeling of eq. 1.
In recent years, various methods developed for the intrusive and non-intrusive (data-driven) reduced-order modeling of first-order linear systems have been extended to the second-order case eq. 1. In the intrusive realm, notable examples include generalizations of the balanced truncation method [26, 16, 34, 11, 45] and transfer function interpolation or moment matching [5, 6, 8, 47, 7]. Balancing-based methods are of particular interest due to their advantageous properties and their interpretation as truncation of internal dynamics associated with small reachability and observability energies. On the other hand, for data-driven methods, extensions of the interpolatory Loewner framework [32, 39, 22], rational vector-fitting [46], operator inference [41, 40, 19] and the AAA algorithm [1] have been developed. In [21], the authors propose a data-driven reformulation of the balanced truncation ($\mathsf{BT}$) method [27, 28], called quadrature-based balanced truncation ($\mathsf{foQuadBT}$). In practice, $\mathsf{foQuadBT}$ typically produces (approximate) $\mathsf{BT}$ surrogate models that are nearly identical to their intrusive counterparts. However, $\mathsf{foQuadBT}$ computes unstructured first-order $\mathsf{BT}$ models, and so a tailored extension suitable for the modeling of second-order systems is needed.
In this work, we propose an extension of the $\mathsf{foQuadBT}$ framework from [21] for the data-driven balanced truncation of second-order dynamical systems eq. 1. Specifically, our method is a data-driven reformulation of the position-velocity balanced truncation ($\mathsf{sopvBT}$) for second-order systems eq. 1 proposed by Reis and Stykel [34]; see Section 2 for details. The major contributions are described as follows.
  • In Section 3, we introduce the structured data-driven balanced truncation framework that we propose, second-order quadrature-based position-velocity balanced truncation ($\mathsf{soQuadpvBT}$). The framework builds on Theorem 1, which derives the intrusive reduced-order quantities in $\mathsf{sopvBT}$ from frequency-response data. These data are evaluations of the transfer function eq. 2 and its position- and velocity-output subsystems. The proposed method applies to any system that satisfies a generalized proportional damping hypothesis, where $\boldsymbol{D}$ is expressed as a linear combination of $\boldsymbol{M}$ and $\boldsymbol{K}$ with potentially frequency-dependent coefficients $f(s)$ and $g(s)$. Parts of the theoretical results presented in this section were derived in the course of writing the dissertation of the corresponding author [36].
  • In Section 4, we derive theoretical results regarding $\mathsf{soQuadpvBT}$. Theorem 3 shows that $\mathsf{soQuadpvBT}$ preserves asymptotic stability when the data are generated by a second-order system eq. 1 with symmetric positive definite ($SPD$) mass, damping and stiffness matrices, and Theorem 4 derives bounds that control the error in the reduced-order model matrices in terms of the associated quadrature error.
  • Because the damping coefficients $f(s)$ and $g(s)$ are typically not known a priori, in Section 5, we propose a computational procedure to estimate these weights from frequency response data in the special cases of Rayleigh and structural damping. Specifically, Theorem 5 provides the gradients needed to infer suitable damping parameters from data.
Several numerical examples are presented in Section 6, which validate the proposed data-driven modeling method and demonstrate its effectiveness. The work is concluded in Section 7.
2 Mathematical background and preliminaries
In this section, we introduce the general damping model that we consider in this paper, and we remind the reader of the essential details of balanced truncation ($\mathsf{BT}$) model reduction for first and second-order systems.
2.1 Generalized proportional damping model
Typically, the mass and stiffness matrices in eq. 1 are assumed to be constant, while the modeling of damping, i.e., the dissipation and conservation of energy in the system, can be significantly more complex. In particular, mathematical models arising in applications such as vibro-acoustics that involve structural dynamics, acoustic wave propagation, or frequency-dependent material properties, only exist in the frequency domain. These models are described by systems of frequency-dependent algebraic equations of the form
$\mathcal{G}^{\operatorname{s}}:\left\{\begin{aligned} \left(s^{2}\boldsymbol{M}+s\boldsymbol{D}(s)+\boldsymbol{K}\right)\boldsymbol{X}(s)&=\boldsymbol{B}_{\operatorname{u}}\boldsymbol{U}(s),\\ \boldsymbol{Y}(s)&=\boldsymbol{C}_{\operatorname{p}}\boldsymbol{X}(s)+\boldsymbol{C}_{\operatorname{v}}\boldsymbol{X}(s).\end{aligned}\right.$ (4)
Here, $\boldsymbol{X}\colon\mathbb{C}\to\mathbb{C}^{n}$, $\boldsymbol{U}\colon\mathbb{C}\to\mathbb{C}^{m}$ and $\boldsymbol{Y}\colon\mathbb{C}\to\mathbb{C}^{p}$ are the Laplace transforms of the state, input, and output vectors from eq. 1, respectively, and the internal damping $\boldsymbol{D}\colon\mathbb{C}\to\mathbb{C}^{n\times n}$ is frequency dependent. See, for example, [4, Sec. 2] and [30] for a more detailed discussion of the different types of internal damping that may appear in the modeling of second-order dynamical systems. We note that in the cases where $\boldsymbol{D}$ is constant, the time- and frequency-domain systems $\mathcal{G}$ in eq. 1 and $\mathcal{G}^{\operatorname{s}}$ in eq. 4 are in fact equivalent formulations related to each other via the Laplace transform.
In this work, we allow the damping term to be frequency-dependent and consider the following generalized form of the proportional damping model
$\boldsymbol{D}(s)=f(s)\boldsymbol{M}+g(s)\boldsymbol{K}.$ (5)
Thus, we require $\boldsymbol{D}$ to be a linear combination of the mass $\boldsymbol{M}$ and stiffness $\boldsymbol{K}$ matrices, where the coefficients $f\colon\mathbb{C}\to\mathbb{C}$ and $g\colon\mathbb{C}\to\mathbb{C}$ are scalar complex-valued frequency-dependent functions. This generalized framework includes, in particular, the following two classical damping models.
Rayleigh damping
Also known as proportional damping, the Rayleigh damping model is given as a constant linear combination of mass and stiffness matrices
$\boldsymbol{D}(s)=\alpha\boldsymbol{M}+\beta\boldsymbol{K},$ (6)
with $\alpha,\beta\geq 0$. This is covered by the general framework eq. 5 by setting $f(s)=\alpha\geq 0$ and $g(s)=\beta\geq 0$. The damping parameters $\alpha$ and $\beta$ allow one to choose the frequency range in which damping is applied, as well as the intensity of the damping.
Structural damping
Also known as hysteretic damping, the structural damping model describes a constant damping effect over the full frequency range using
$\boldsymbol{D}(s)=\mathfrak{i}\frac{\eta}{s}\boldsymbol{K},$ (7)
where $\eta\geq 0$ is a material-dependent structural loss factor. In the general framework eq. 5, this damping model is given by setting $f(s)=0$ and $g(s)=\mathfrak{i}\tfrac{\eta}{s}$.
2.2 Balanced truncation model reduction
Consider first-order linear systems of the form
$\mathcal{G}_{\operatorname{fo}}:\left\{\begin{aligned} \boldsymbol{E}\dot{\boldsymbol{q}}(t)&=\boldsymbol{A}\boldsymbol{q}(t)+\boldsymbol{B}\boldsymbol{u}(t),\quad\boldsymbol{q}(0)=\boldsymbol{0},\\ \boldsymbol{y}(t)&=\boldsymbol{C}\boldsymbol{q}(t),\end{aligned}\right.$ (8)
where $\boldsymbol{E},\boldsymbol{A}\in\mathbb{R}^{n_{\operatorname{fo}}\times n_{\operatorname{fo}}}$, $\boldsymbol{B}\in\mathbb{R}^{n_{\operatorname{fo}}\times m}$, $\boldsymbol{C}\in\mathbb{R}^{p\times n_{\operatorname{fo}}}$ and $\boldsymbol{q}\colon[0,\infty)\to\mathbb{R}^{n_{\operatorname{fo}}}$. We assume that $\boldsymbol{E}$ is nonsingular, that the realization of the system eq. 8 is controllable and observable, and that the system eq. 8 is asymptotically stable, i.e., that the eigenvalues of the matrix pencil $\lambda\boldsymbol{E}-\boldsymbol{A}$ lie in the open left half of the complex plane. The core idea of $\mathsf{BT}$ is the simultaneous diagonalization of two symmetric positive semi-definite matrices, called system Gramians, and the truncation of states corresponding to small eigenvalues of these Gramians [28, 27]. The system Gramians $\boldsymbol{P},\boldsymbol{E}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}\boldsymbol{E}\in\mathbb{R}^{n_{\operatorname{fo}}\times n_{\operatorname{fo}}}$, also known as the controllability and observability Gramians, respectively, are defined via
$\boldsymbol{P}$ $=\frac{1}{2\pi}\int\limits_{-\infty}^{\infty}(\mathfrak{i}z\boldsymbol{E}-\boldsymbol{A})^{-1}\boldsymbol{B}\boldsymbol{B}^{\mkern-1.5mu\mathsf{T}}(\mathfrak{i}z\boldsymbol{E}-\boldsymbol{A})^{-\mathsf{H}}\operatorname{d}z,\quad\text{and}$ (9a)
$\boldsymbol{Q}$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(\mathfrak{i}z\boldsymbol{E}-\boldsymbol{A})^{-\mathsf{H}}\boldsymbol{C}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{C}(\mathfrak{i}z\boldsymbol{E}-\boldsymbol{A})^{-1}\operatorname{d}z.$ (9b)
Under the assumptions stated above, the matrices $\boldsymbol{P}$ and $\boldsymbol{Q}$ are $SPD$, and thus admit Cholesky factorizations of the form $\boldsymbol{P}=\boldsymbol{R}\boldsymbol{R}^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{Q}=\boldsymbol{L}\boldsymbol{L}^{\mkern-1.5mu\mathsf{T}}$ where $\boldsymbol{R},\boldsymbol{L}\in\mathbb{R}^{n_{\operatorname{fo}}\times n_{\operatorname{fo}}}$. The system-theoretic significance of these matrices is that $\boldsymbol{P}$ and $\boldsymbol{E}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}\boldsymbol{E}$ respectively quantify the controllability and observability of a state $\boldsymbol{q}$. Then, $\mathsf{BT}$ computes a reduced-order model by removing components of the state space that correspond to small controllability and observability energies in the balanced basis. The appeal of $\mathsf{BT}$ comes from the preservation of asymptotic stability in the reduced-order model [31] and the associated a priori error bound on the $\mathcal{H}_{\infty}$ norm [18]. For further details on balancing-related model reduction for first-order linear systems eq. 8, we refer the reader to [2, 23, 9] and references therein.
2.3 Balanced truncation for second-order systems
There have been several attempts to generalize the ideas of classical $\mathsf{BT}$ to second-order systems [26, 34, 16, 45]. The unifying feature of the proposed methods is the reformulation of the second-order system in eq. 1 as an equivalent $2n$-dimensional first-order system eq. 8. This can always be accomplished by introducing the first-order state vector $\boldsymbol{q}^{\mkern-1.5mu\mathsf{T}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\begin{bmatrix}\boldsymbol{x}^{\mkern-1.5mu\mathsf{T}}&\dot{\boldsymbol{x}}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix}$ and re-organizing the second-order dynamics in eq. 1 accordingly. The resulting $2n$-dimensional system eq. 8 is given in the so-called first companion form, with the system matrices
$\boldsymbol{E}=\begin{bmatrix}\boldsymbol{I}_{n}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{M}\end{bmatrix},\quad\boldsymbol{A}=\begin{bmatrix}\phantom{-}\boldsymbol{0}&\phantom{-}\boldsymbol{I}_{n}\\ -\boldsymbol{K}&-\boldsymbol{D}\end{bmatrix},\quad\boldsymbol{C}=\begin{bmatrix}\boldsymbol{C}_{\operatorname{p}}&\boldsymbol{C}_{\operatorname{v}}\end{bmatrix},\quad\boldsymbol{B}=\begin{bmatrix}\boldsymbol{0}\\ \boldsymbol{B}_{\operatorname{u}}\end{bmatrix},$ (10)
where $\boldsymbol{I}_{n}$ denotes the $n\times n$ identity matrix. As in Section 2.2, we assume throughout this paper that the mass matrix $\boldsymbol{M}$ in eq. 1 is nonsingular, that the realization of the system eq. 1 is controllable and observable, and that the system eq. 1 is asymptotically stable, i.e., the eigenvalues of the matrix pencil $\lambda^{2}\boldsymbol{M}+\lambda\boldsymbol{D}+\boldsymbol{K}$ lie in the open left half of the complex plane. Following these assumptions about the second-order system eq. 1, the corresponding first-order realization eq. 10 has a nonsingular $\boldsymbol{E}$ matrix, and is asymptotically stable as well as controllable and observable [24].
Let the first-order system Gramians in eq. 9 be partitioned according to the block structure in the companion form eq. 10 as
$\boldsymbol{P}=\begin{bmatrix}\boldsymbol{P}_{\operatorname{p}}&\boldsymbol{P}_{12}\\[2.0pt] \boldsymbol{P}_{12}^{\mkern-1.5mu\mathsf{T}}&\boldsymbol{P}_{\operatorname{v}}\end{bmatrix}\quad\text{and}\quad\boldsymbol{E}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}\boldsymbol{E}=\begin{bmatrix}\boldsymbol{Q}_{\operatorname{p}}&\boldsymbol{Q}_{12}\boldsymbol{M}\\[2.0pt] \boldsymbol{M}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}_{12}^{\mkern-1.5mu\mathsf{T}}&\boldsymbol{M}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}_{\operatorname{v}}\boldsymbol{M}\end{bmatrix}.$ (11)
The submatrices $\boldsymbol{P}_{\operatorname{p}},\boldsymbol{Q}_{\operatorname{p}}\in\mathbb{R}^{n\times n}$ are the so-called position-controllability and position-observability Gramians, while $\boldsymbol{P}_{\operatorname{v}},\boldsymbol{M}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}_{\operatorname{v}}\boldsymbol{M}\in\mathbb{R}^{n\times n}$ are the velocity-controllability and velocity-observability Gramians. Because $\boldsymbol{P}$ and $\boldsymbol{Q}$ are SPD, so too are their diagonal submatrices. The Gramians $\boldsymbol{P}_{\operatorname{p}}$ and $\boldsymbol{M}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}_{\operatorname{v}}\boldsymbol{M}$ can alternatively be expressed via contour integrals in terms of the input-to-state and state-to-output mappings of the system eq. 1 in the frequency domain
$\boldsymbol{P}_{\operatorname{p}}$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}\left(-z^{2}\boldsymbol{M}+\mathfrak{i}z\boldsymbol{D}+\boldsymbol{K}\right)^{-1}\boldsymbol{B}_{\operatorname{u}}\boldsymbol{B}_{\operatorname{u}}^{\mkern-1.5mu\mathsf{T}}\left(-z^{2}\boldsymbol{M}+\mathfrak{i}z\boldsymbol{D}+\boldsymbol{K}\right)^{-\mathsf{H}}\operatorname{d}z,$ (12a)
$\boldsymbol{Q}_{\operatorname{v}}$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}\left(-z^{2}\boldsymbol{M}+\mathfrak{i}z\boldsymbol{D}+\boldsymbol{K}\right)^{-\mathsf{H}}(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}z\boldsymbol{C}_{\operatorname{v}})^{\mathsf{H}}$ (12b)
[…]
see, e.g., [33, Prop. 2.1] and [15, Sec. 4.4]. The methods proposed in [26, 34] correspond to balancing selected combinations of the second-order controllability and observability Gramians in eq. 11 with subsequent truncation. In particular, the so-called free-velocity balanced truncation from [26] and the position-velocity balanced truncation ($\mathsf{sopvBT}$) from [34] correspond to the simultaneous diagonalization of $\boldsymbol{P}_{\operatorname{p}},\boldsymbol{Q}_{\operatorname{p}}$ and $\boldsymbol{P}_{\operatorname{p}},\boldsymbol{M}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}_{\operatorname{v}}\boldsymbol{M}$, respectively. Algorithmically, balancing and truncation are achieved simultaneously by using an appropriate generalization of the square-root algorithm from the first-order system case [25, 42]. For a more detailed overview of balanced truncation for second-order systems and all the different variations, we refer the reader to [45, 37, 11].
The primary focus of this work lies on the $\mathsf{sopvBT}$ method from [34]. The main steps of the approach are summarized in Algorithm 1. In the case of symmetric second-order systems, i.e., the mass, damping, and stiffness matrices in eq. 1 are $SPD$ with $\boldsymbol{B}_{\operatorname{u}}=\boldsymbol{C}_{\operatorname{p}}^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{0}$, asymptotic stability is preserved by $\mathsf{sopvBT}$. In fact, in this particular situation, we have that $\boldsymbol{P}_{\operatorname{p}}=\boldsymbol{Q}_{\operatorname{v}}$; see [34, Thm. 3.1]. Note also that Algorithm 1 can be applied if either the position or velocity-output terms in eq. 1 are zero.
1
Input: Second-order system matrices $\boldsymbol{M},\boldsymbol{D},\boldsymbol{K},\boldsymbol{B}_{\operatorname{u}},\boldsymbol{C}_{\operatorname{p}},\boldsymbol{C}_{\operatorname{v}}$, reduction order $r$.
Output: Reduced-order system matrices ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{D}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}},{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}},{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ in eq. 3.
2
3Compute Cholesky factorizations $\boldsymbol{P}=\boldsymbol{R}\boldsymbol{R}^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{Q}=\boldsymbol{L}\boldsymbol{L}^{\mkern-1.5mu\mathsf{T}}$ of the system Gramians eq. 9 using the companion form eq. 10, where the factors are partitioned as
$\boldsymbol{R}=\begin{bmatrix}\boldsymbol{R}_{\operatorname{p}}\\ \star\end{bmatrix}\quad\text{and}\quad\boldsymbol{L}=\begin{bmatrix}\star\\ \boldsymbol{L}_{\operatorname{v}}\end{bmatrix},$
with $\boldsymbol{R}_{\operatorname{p}},\boldsymbol{L}_{\operatorname{v}}\in\mathbb{R}^{n\times n}$; the $\star$ marks submatrices that are not needed.
4
5Compute the singular value decomposition
$\boldsymbol{L}_{\operatorname{v}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{M}\boldsymbol{R}_{\operatorname{p}}=\begin{bmatrix}\boldsymbol{U}_{1}&\boldsymbol{U}_{2}\end{bmatrix}\begin{bmatrix}\boldsymbol{S}_{1}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{S}_{2}\end{bmatrix}\begin{bmatrix}\boldsymbol{V}_{1}^{\mkern-1.5mu\mathsf{T}}\\[2.0pt] \boldsymbol{V}_{2}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},$
where $\boldsymbol{U}_{1},\boldsymbol{V}_{1}\in\mathbb{R}^{n\times r}$ have orthonormal columns and $\boldsymbol{S}_{1}\in\mathbb{R}^{r\times r}$ is diagonal containing the $r$ largest singular values.
6
7Compute the reduced-order matrices by projection
${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}$ $=\boldsymbol{W}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{M}\boldsymbol{T}$ $=\boldsymbol{I}_{r},$ (13a)
${\widetilde{\boldsymbol{D}\mkern 2.0mu}\mkern-2.0mu}{}$ $=\boldsymbol{W}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{D}\boldsymbol{T}$ $=\boldsymbol{S}_{1}^{-1/2}\boldsymbol{U}_{1}^{\mkern-1.5mu\mathsf{T}}\left(\boldsymbol{L}_{\operatorname{v}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{D}\boldsymbol{R}_{\operatorname{p}}\right)\boldsymbol{V}_{1}\boldsymbol{S}_{1}^{-1/2},$ (13b)
${\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}$ $=\boldsymbol{W}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{K}\boldsymbol{T}$ $=\boldsymbol{S}_{1}^{-1/2}\boldsymbol{U}_{1}^{\mkern-1.5mu\mathsf{T}}\left(\boldsymbol{L}_{\operatorname{v}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{K}\boldsymbol{R}_{\operatorname{p}}\right)\boldsymbol{V}_{1}\boldsymbol{S}_{1}^{-1/2},$ (13c)
${\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}$ $=\boldsymbol{W}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{B}_{\operatorname{u}}$ $=\boldsymbol{S}_{1}^{-1/2}\boldsymbol{U}_{1}^{\mkern-1.5mu\mathsf{T}}\left(\boldsymbol{L}_{\operatorname{v}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{B}_{\operatorname{u}}\right),$ (13d)
${\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}}$ $=\boldsymbol{C}_{\operatorname{p}}\boldsymbol{T}$ $=\left(\boldsymbol{C}_{\operatorname{p}}\boldsymbol{R}_{\operatorname{p}}\right)\boldsymbol{V}_{1}\boldsymbol{S}_{1}^{-1/2},$ (13e)
${\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ $=\boldsymbol{C}_{\operatorname{v}}\boldsymbol{T}$ $=\left(\boldsymbol{C}_{\operatorname{v}}\boldsymbol{R}_{\operatorname{p}}\right)\boldsymbol{V}_{1}\boldsymbol{S}_{1}^{-1/2}.$ (13f)
Algorithm 1 Second-order position-velocity $\mathsf{BT}$ ($\mathsf{sopvBT}$).
Remark 1.
In the more complex case of frequency-dependent damping eq. 4, one can still formulate structured Gramians via the integral representations in eq. 12 by replacing the constant $\boldsymbol{D}$ with the frequency-dependent $\boldsymbol{D}(s)$. Then, if the integrals converge, the structured Gramians can be approximated by numerical quadrature rules, and Algorithm 1 can be applied; see, for example, the work on structure-preserving model reduction of integral-differential equations in [15].
Remark 2.
The reason only $\mathsf{sopvBT}$ is considered here is due to the integral representations of the second-order system Gramians eq. 12. These representations are given in terms of the input-to-state and state-to-output maps of the second-order system eq. 1, and they will allow us to facilitate a data-driven reformulation of the method expressed in terms of transfer function data eq. 2. Even though one can also write down integral formulations of $\boldsymbol{P}_{\operatorname{v}}$ and $\boldsymbol{Q}_{\operatorname{p}}$, the structure of these integrals is not as amenable to the tools used in Section 3. In particular, it is not clear what the “data” needed in the computations correspond to.
3 Data-driven second-order balanced truncation
The $\mathsf{sopvBT}$ method in Algorithm 1 is intrusive in the sense that it requires an explicit state-space model of the second-order system eq. 1 to compute the Cholesky factors $\boldsymbol{R}_{\operatorname{p}},\boldsymbol{L}_{\operatorname{v}}\in\mathbb{R}^{n\times n}$ and to compute the reduced-order model eq. 3 by projecting the full-order model matrices eq. 13. In this section, we present a data-driven reformulation of $\mathsf{sopvBT}$. Our only assumption is that we can sample the full-order transfer function eq. 2 in a prescribed range of frequencies along the imaginary axis $\mathfrak{i}\mathbb{R}$ as well as the coefficient functions in eq. 5, which we assume for now are known a priori. The resulting method, which we call second-order quadrature-based position-velocity balanced truncation ($\mathsf{soQuadpvBT}$), is a generalization of the $\mathsf{foQuadBT}$ framework for first-order systems from [21]. Our method allows for the computation of (approximate) $\mathsf{sopvBT}$-based second-order surrogate models eq. 3 directly from frequency-response data.
We begin with the following observation: The matrices $\boldsymbol{U}_{1}$, $\boldsymbol{Y}_{1}$ and $\boldsymbol{\Sigma}_{1}$ in eq. 13 stem from the truncated singular value decomposition of $\boldsymbol{L}_{\operatorname{v}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{M}\boldsymbol{R}_{\operatorname{p}}$. Under the assumption that $\boldsymbol{D}(s)=f(s)\boldsymbol{M}+g(s)\boldsymbol{K}$, it follows that the intrusive $\mathsf{sopvBT}$ reduced-order model constructed via Algorithm 1 is fully characterized by the five quantities
$\boldsymbol{L}_{\operatorname{v}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{M}\boldsymbol{R}_{\operatorname{p}},\quad\boldsymbol{L}_{\operatorname{v}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{K}\boldsymbol{R}_{\operatorname{p}},\quad\boldsymbol{L}_{\operatorname{v}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{B}_{\operatorname{u}},\quad\boldsymbol{C}_{\operatorname{p}}\boldsymbol{R}_{\operatorname{p}}\quad\text{and}\quad\boldsymbol{C}_{\operatorname{v}}\boldsymbol{R}_{\operatorname{p}}.$ (14)
Along with the integral formulations of the position-reachability and velocity-observability Gramians $\boldsymbol{P}_{\operatorname{p}}\in\mathbb{R}^{n\times n}$ and $\boldsymbol{M}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}_{\operatorname{v}}\boldsymbol{M}\in\mathbb{R}^{n\times n}$ in eq. 12, this observation suggests a natural extension of $\mathsf{foQuadBT}$ for the second-order system eq. 1. In [21], the authors arrive at a fully data-driven reformulation of the first-order $\mathsf{BT}$ method by replacing the exact square-root factors of $\boldsymbol{R}$ and $\boldsymbol{L}$ with low-rank factors derived from numerical quadrature rules applied to the integral representations of the Gramians eq. 9. The quadrature-based factors are never explicitly formed and the method only makes reference to transfer function evaluations at the underlying quadrature nodes. We use this idea as a blueprint to develop a data-driven reformulation of $\mathsf{sopvBT}$. Specifically, we use implicit numerical quadrature rules to derive low-rank approximations of the exact factors $\boldsymbol{R}_{\operatorname{p}}$ and $\boldsymbol{L}_{\operatorname{v}}$, and ultimately show how to realize the quadrature-based approximations of the intrusive reduced-order quantities in eq. 14 non-intrusively using only transfer function data.
Recall the integral representations of the Gramians eq. 12, but with the frequency-dependent $\boldsymbol{D}(s)$ in place of $\boldsymbol{D}$. For compactness of the presentation, we introduce the notation
$\boldsymbol{\varphi}\colon\mathbb{C}\to\mathbb{C}^{n\times n}\quad\text{with}\quad\boldsymbol{\varphi}(s)\stackrel{{\scriptstyle\operatorname{def}}}{{=}}s^{2}\boldsymbol{M}+s\boldsymbol{D}(s)+\boldsymbol{K}$ (15)
for the resolvent term of the second-order transfer function eq. 2. Consider a numerical quadrature rule defined by the nodes $\mathfrak{i}\zeta_{1},\ldots,\mathfrak{i}\zeta_{J}\in\mathfrak{i}\mathbb{R}$ and the corresponding weights $\varrho_{1}^{2},\ldots,\varrho_{J}^{2}\in\mathbb{R}$. Applying this quadrature rule to $\boldsymbol{P}_{\operatorname{p}}$ in eq. 12a reveals the approximate factorization
$\boldsymbol{P}_{\operatorname{p}}\approx\sum\limits_{j=1}^{J}\varrho_{j}^{2}\boldsymbol{\varphi}(\mathfrak{i}\zeta_{j})^{-1}\boldsymbol{B}_{\operatorname{u}}\left(\boldsymbol{\varphi}(\mathfrak{i}\zeta_{j})^{-1}\boldsymbol{B}_{\operatorname{u}}\right)^{\mathsf{H}}={\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}^{\mathsf{H}},$
where the factor ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{n\times mJ}$ is defined as
${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\begin{bmatrix}\varrho_{1}\boldsymbol{\varphi}(\mathfrak{i}\zeta_{1})^{-1}\boldsymbol{B}_{\operatorname{u}}&\varrho_{2}\boldsymbol{\varphi}(\mathfrak{i}\zeta_{2})^{-1}\boldsymbol{B}_{\operatorname{u}}&\ldots&\varrho_{J}\boldsymbol{\varphi}(\mathfrak{i}\zeta_{J})^{-1}\boldsymbol{B}_{\operatorname{u}}\end{bmatrix}.$ (16)
A similar factorization can be derived for $\boldsymbol{Q}_{\operatorname{v}}$ in eq. 12b. Given another numerical quadrature rule defined by the nodes $\mathfrak{i}\vartheta_{1},\ldots,\mathfrak{i}\vartheta_{K}\in\mathfrak{i}\mathbb{R}$ and the corresponding weights $\varphi_{1}^{2},\ldots,\varphi_{K}^{2}\in\mathbb{R}$, applying this quadrature rule then yields
$\boldsymbol{Q}_{\operatorname{v}}\approx\sum\limits_{k=1}^{K}\varphi_{k}^{2}\left(\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{k}\boldsymbol{C}_{\operatorname{v}}\right)\boldsymbol{\varphi}(\mathfrak{i}\vartheta_{k})^{-1}\right)^{\mathsf{H}}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{k}\boldsymbol{C}_{\operatorname{v}}\right)\boldsymbol{\varphi}(\mathfrak{i}\vartheta_{k})^{-1}={\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}},$
where the factor ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}\in\mathbb{C}^{n\times pK}$ is defined according to
${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\begin{bmatrix}\varphi_{1}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{1}\boldsymbol{C}_{\operatorname{v}}\right)\boldsymbol{\varphi}(\mathfrak{i}\vartheta_{1})^{-1}\\ \varphi_{2}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{2}\boldsymbol{C}_{\operatorname{v}}\right)\boldsymbol{\varphi}(\mathfrak{i}\vartheta_{2})^{-1}\\[2.0pt] \vdots\\[2.0pt] \varphi_{K}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{K}\boldsymbol{C}_{\operatorname{v}}\right)\boldsymbol{\varphi}(\mathfrak{i}\vartheta_{K})^{-1}\end{bmatrix}.$ (17)
Replacing the exact Cholesky factors $\boldsymbol{R}_{\operatorname{p}}$ and $\boldsymbol{L}_{\operatorname{v}}$ in Algorithm 1 with the quadrature-based factors ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}$ and ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ from eqs. 16 and 17 in forming the matrices eq. 14 already yields a low-rank implementation of $\mathsf{sopvBT}$. As we show next in Theorem 1, the significant implication of this low-rank formulation is that the quadrature-based approximations to the intrusive reduced-order quantities in eq. 14 can be computed non-intrusively from transfer function samples and evaluations of the damping coefficients $f(s)$ and $g(s)$ in eq. 5.
Before moving on to the results, we introduce the following position and velocity-output transfer functions $\boldsymbol{G}_{\operatorname{p}}\colon\mathbb{C}\to\mathbb{C}^{p\times m}$ and $\boldsymbol{G}_{\operatorname{v}}\colon\mathbb{C}\to\mathbb{C}^{p\times m}$ of the system eq. 1, defined as
$\boldsymbol{G}_{\operatorname{p}}(s)\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{C}_{\operatorname{p}}\boldsymbol{\varphi}(s)^{-1}\boldsymbol{B}_{\operatorname{u}}\quad\text{and}\quad\boldsymbol{G}_{\operatorname{v}}(s)=s\boldsymbol{C}_{\operatorname{v}}\boldsymbol{\varphi}(s)^{-1}\boldsymbol{B}_{\operatorname{u}}.$ (18)
The two functions $\boldsymbol{G}_{\operatorname{p}}$ and $\boldsymbol{G}_{\operatorname{v}}$ are each the transfer function of a second-order system eq. 1 with purely position and velocity outputs, respectively. In general, we have that $\boldsymbol{G}(s)=\boldsymbol{G}_{\operatorname{p}}(s)+s\boldsymbol{G}_{\operatorname{v}}(s)$. For ease of presentation, we introduce the following notation for indexing the entries of block matrices: Given a block matrix $\boldsymbol{X}\in\mathbb{C}^{pK\times mJ}$, we use
$\boldsymbol{X}_{\boldsymbol{j},\boldsymbol{k}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{X}_{(k-1)p+1:kp,(j-1)m+1:jm}\in\mathbb{C}^{K\times J},$
to denote the block submatrix of $\boldsymbol{X}$ containing the rows $(k-1)p+1,\ldots,kp$ and the columns $(j-1)m+1,\ldots,jm$. If $p=1$ (or $m=1$), we instead write $\boldsymbol{X}_{:,\boldsymbol{j}}$ (or $\boldsymbol{X}_{\boldsymbol{k},:}$) to denote the $j$-th (or $k$-th) block entry of $\boldsymbol{X}$.
To simplify the proof of the main result Theorem 1, we first show the following lemma from [17].
Lemma 1.
For matrices $\boldsymbol{X},\boldsymbol{Y}\in\mathbb{C}^{n\times n}$, the following two identities
$(s\boldsymbol{X}+\boldsymbol{Y})^{-1}\boldsymbol{X}(z\boldsymbol{X}+\boldsymbol{Y})^{-1}$ $=\frac{(z\boldsymbol{X}+\boldsymbol{Y})^{-1}-(s\boldsymbol{X}+\boldsymbol{Y})^{-1}}{s-z},$ (19a)
$(s\boldsymbol{X}+\boldsymbol{Y})^{-1}\boldsymbol{Y}(z\boldsymbol{X}+\boldsymbol{Y})^{-1}$ $=\frac{z(z\boldsymbol{X}+\boldsymbol{Y})^{-1}-s(s\boldsymbol{X}+\boldsymbol{Y})^{-1}}{z-s}$ (19b)
hold for any $s,z\in\mathbb{C}$, with $s\neq z$ and such that $s\boldsymbol{X}+\boldsymbol{Y}$ and $z\boldsymbol{X}+\boldsymbol{Y}$ are nonsingular.
Proof.
For any $s,z\in\mathbb{C}$ for which $s\boldsymbol{X}+\boldsymbol{Y}$ and $z\boldsymbol{X}+\boldsymbol{Y}$ are nonsingular, we observe that
$(s-z)(s\boldsymbol{X}+\boldsymbol{Y})^{-1}\boldsymbol{X}(z\boldsymbol{X}+\boldsymbol{Y})^{-1}$
$\leavevmode\nobreak\ =(s\boldsymbol{X}+\boldsymbol{Y})^{-1}\big{(}(s\boldsymbol{X}+\boldsymbol{Y})-(z\boldsymbol{X}+\boldsymbol{Y})\big{)}(z\boldsymbol{X}+\boldsymbol{Y})^{-1}$
$\leavevmode\nobreak\ =(z\boldsymbol{X}+\boldsymbol{Y})^{-1}-(s\boldsymbol{X}+\boldsymbol{Y})^{-1}$
holds, thus proving eq. 19a. Similarly, we observe that
$(z-s)(s\boldsymbol{X}+\boldsymbol{Y})^{-1}\boldsymbol{Y}(z\boldsymbol{X}+\boldsymbol{Y})^{-1}$
$\leavevmode\nobreak\ =(s\boldsymbol{X}+\boldsymbol{Y})^{-1}\big{(}(sz\boldsymbol{X}+z\boldsymbol{Y})-(sz\boldsymbol{X}+z\boldsymbol{Y})\big{)}(z\boldsymbol{X}+\boldsymbol{Y})^{-1}$
$\leavevmode\nobreak\ =z(z\boldsymbol{X}+\boldsymbol{Y})^{-1}-s(s\boldsymbol{X}+\boldsymbol{Y})^{-1},$
holds, which proves eq. 19b. ∎
Theorem 1.
Define the functions $d\colon\mathbb{C}\to\mathbb{C}$, $n\colon\mathbb{C}\to\mathbb{C}$ and $h\colon\mathbb{C}\to\mathbb{C}$ as
$d(s)\stackrel{{\scriptstyle\operatorname{def}}}{{=}}1+sg(s),\quad n(s)\stackrel{{\scriptstyle\operatorname{def}}}{{=}}s^{2}+sf(s)\quad\text{and}\quad h(s)\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\frac{n(s)}{d(s)},$ (20)
where $f(s)$ and $g(s)$ are the damping coefficient functions from eq. 5. Suppose the left and right quadrature nodes $\mathfrak{i}\vartheta_{1},\ldots,\mathfrak{i}\vartheta_{K}$ and $\mathfrak{i}\zeta_{1},\ldots,\mathfrak{i}\zeta_{J}$ in eqs. 16 and 17 are such that
$\zeta_{j}\neq 0,\quad d(\mathfrak{i}\vartheta_{k})\neq 0,\quad d(\mathfrak{i}\zeta_{j})\neq 0,\quad h(\mathfrak{i}\vartheta_{k})\neq h(\mathfrak{i}\zeta_{j}),$
and so that the matrices $\boldsymbol{\varphi}(\mathfrak{i}\zeta_{j})$ and $\boldsymbol{\varphi}(\mathfrak{i}\vartheta_{k})$ are invertible for all $j=1,\ldots,J$ and $k=1,\ldots,K$, where $\boldsymbol{\varphi}$ is defined in eq. 15. Let the quadrature-based factors ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}$ and ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ be given as in eqs. 16 and 17, and define the matrices
$\mathbb{M}$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{M}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{pK\times mJ},\qquad$ $\mathbb{K}$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{K}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{pK\times mJ},$ (21a)
$\mathbb{B}_{\operatorname{u}}$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{B}_{\operatorname{u}}\in\mathbb{C}^{pK\times m},\qquad$ $\mathbb{C}_{\operatorname{p}}$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{C}_{\operatorname{p}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{p\times mJ},$ (21b)
$\mathbb{C}_{\operatorname{v}}$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{C}_{\operatorname{v}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{p\times mJ}.$ (21c)
Then, the matrices eq. 21 are given block-entrywise by
$\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}$ $=-\frac{\varphi_{k}\varrho_{j}}{d(\mathfrak{i}\vartheta_{k})d(\mathfrak{i}\zeta_{j})}\frac{d(\mathfrak{i}\vartheta_{k})\boldsymbol{G}(\mathfrak{i}\vartheta_{k})-d(\mathfrak{i}\zeta_{j})\left(\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j})+\frac{\vartheta_{k}}{\zeta_{j}}\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j})\right)}{h(\mathfrak{i}\vartheta_{k})-h(\mathfrak{i}\zeta_{j})},$ (22a)
$\mathbb{K}_{\boldsymbol{k},\boldsymbol{j}}$ $=\hphantom{-}\frac{\varphi_{k}\varrho_{j}}{d(\mathfrak{i}\vartheta_{k})d(\mathfrak{i}\zeta_{j})}\frac{n(\mathfrak{i}\vartheta_{k})\boldsymbol{G}(\mathfrak{i}\vartheta_{k})-n(\mathfrak{i}\zeta_{j})\left(\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j})+\frac{\vartheta_{k}}{\zeta_{j}}\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j})\right)}{h(\mathfrak{i}\vartheta_{k})-h(\mathfrak{i}\zeta_{j})},$ (22b)
$\left(\mathbb{B}_{\operatorname{u}}\right)_{\boldsymbol{k},:}$ $=\hphantom{-}\varphi_{k}\boldsymbol{G}(\mathfrak{i}\vartheta_{k}),\quad\left(\mathbb{C}_{\operatorname{p}}\right)_{:,\boldsymbol{j}}=\varrho_{j}\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j})\quad\text{and}\quad\left(\mathbb{C}_{\operatorname{v}}\right)_{:,\boldsymbol{j}}=\frac{\varrho_{j}}{\mathfrak{i}\zeta_{j}}\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j}),$ (22c)
where $\boldsymbol{G}_{\operatorname{p}}$ and $\boldsymbol{G}_{\operatorname{v}}$ are defined as in eq. 18.
Proof.
For any indices $i,\ell\in\mathbb{N}$, we introduce the matrix
$\boldsymbol{I}_{i,\ell}=\begin{bmatrix}\boldsymbol{e}_{(i-1)\ell+1}&\boldsymbol{e}_{(i-1)\ell+2}&\cdots&\boldsymbol{e}_{i\ell}\end{bmatrix}\in\mathbb{R}^{n\times\ell},$ (23)
where $\boldsymbol{e}_{i}\in\mathbb{R}^{n}$ is the $i$-th canonical basis vector with $1$ at the $i$-th position and 00 everywhere else. Note that $\boldsymbol{I}_{i,\ell}$ contains a subset of the columns of the $n\times n$ identity and has the effect of extracting the columns $(i-1)\ell+1$ through $i\ell$ of a matrix by right multiplication. The formulae for $\mathbb{B}_{\operatorname{u}}$, $\mathbb{C}_{\operatorname{p}}$ and $\mathbb{C}_{\operatorname{v}}$ in eq. 22c are a direct consequence of their construction in eq. 21 and the definitions of the quadrature-based factors ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}$ and ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$. We easily observe that
[…]
holds, where $\boldsymbol{I}_{k,p}$ is defined according to eq. 23. Similarly, we have that
$\left(\mathbb{C}_{\operatorname{p}}\right)_{:,\boldsymbol{j}}$ $=\boldsymbol{C}_{\operatorname{p}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\boldsymbol{I}_{j,m}=\varrho_{j}\boldsymbol{C}_{\operatorname{p}}\boldsymbol{\varphi}(\mathfrak{i}\zeta_{j})^{-1}\boldsymbol{B}_{\operatorname{u}}=\varrho_{j}\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j})\quad\text{and}$
$\left(\mathbb{C}_{\operatorname{v}}\right)_{:,\boldsymbol{j}}$ $=\boldsymbol{C}_{\operatorname{v}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\boldsymbol{I}_{j,m}=\varrho_{j}\boldsymbol{C}_{\operatorname{v}}\boldsymbol{\varphi}(\mathfrak{i}\zeta_{j})^{-1}\boldsymbol{B}_{\operatorname{u}}=\frac{\varrho_{j}}{\mathfrak{i}\zeta_{j}}\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j}).$
For the remaining formulae eqs. 22a and 22b, we first note that so long as $d(s)\neq 0$, it holds that
$\boldsymbol{\varphi}(s)^{-1}$ $=\big{(}s^{2}\boldsymbol{M}+s\boldsymbol{D}(s)+\boldsymbol{K}\big{)}^{-1}$
$=\Big{(}\big{(}s^{2}+sf(s)\big{)}\boldsymbol{M}+\big{(}1+sg(s)\big{)}\boldsymbol{K}\Big{)}^{-1}$
$=\frac{1}{1+sg(s)}\left(\frac{s^{2}+sf(s)}{1+sg(s)}\boldsymbol{M}+\boldsymbol{K}\right)^{-1}$
$=\frac{1}{d(s)}\big{(}h(s)\boldsymbol{M}+\boldsymbol{K}\big{)}^{-1}$
$=\frac{1}{d(s)}\boldsymbol{\phi}\left(h(s)\right)^{-1},$
where $\boldsymbol{\phi}(s)\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\left(s\boldsymbol{M}+\boldsymbol{K}\right)$. To prove eqs. 22a and 22b, we use the resolvent identities from Lemma 1. Under the assumption that $h(\mathfrak{i}\vartheta_{k})\neq h(\mathfrak{i}\zeta_{j})$ for all $k$ and $j$, applying eq. 19a reveals that
$\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}$ $=\boldsymbol{I}_{k,p}^{\mkern-1.5mu\mathsf{T}}\left({\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{M}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\right)\boldsymbol{I}_{j,m}$
$=\varphi_{k}\varrho_{j}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{k}\boldsymbol{C}_{\operatorname{v}}\right)\boldsymbol{\varphi}(\mathfrak{i}\vartheta_{k})^{-1}\boldsymbol{M}\boldsymbol{\varphi}(\mathfrak{i}\zeta_{j})^{-1}\boldsymbol{B}_{\operatorname{u}}$
$=\varphi_{k}\varrho_{j}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{k}\boldsymbol{C}_{\operatorname{v}}\right)\frac{\boldsymbol{\phi}\left(h(\mathfrak{i}\vartheta_{k})\right)^{-1}\boldsymbol{M}\boldsymbol{\phi}\left(h(\mathfrak{i}\zeta_{j})\right)^{-1}}{d(\mathfrak{i}\vartheta_{k})d(\mathfrak{i}\zeta_{j})}\boldsymbol{B}_{\operatorname{u}}$
$=-\varphi_{k}\varrho_{j}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{k}\boldsymbol{C}_{\operatorname{v}}\right)\frac{\boldsymbol{\phi}\left(h(\mathfrak{i}\vartheta_{k})\right)^{-1}-\boldsymbol{\phi}\left(h(\mathfrak{i}\zeta_{j})\right)^{-1}}{d(\mathfrak{i}\vartheta_{k})d(\mathfrak{i}\zeta_{j})\left(h(\mathfrak{i}\vartheta_{k})-h(\mathfrak{i}\zeta_{j})\right)}\boldsymbol{B}_{\operatorname{u}}$
$=-\frac{\varphi_{k}\varrho_{j}}{d(\mathfrak{i}\vartheta_{k})d(\mathfrak{i}\zeta_{j})}\frac{d(\mathfrak{i}\vartheta_{k})\boldsymbol{G}(\mathfrak{i}\vartheta_{k})-d(\mathfrak{i}\zeta_{j})\left(\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j})+\frac{\vartheta_{k}}{\zeta_{j}}\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j})\right)}{h(\mathfrak{i}\vartheta_{k})-h(\mathfrak{i}\zeta_{j})},$
where the last line follows from the relationship $\boldsymbol{\varphi}(s)^{-1}=d(s)^{-1}\boldsymbol{\phi}\left(h(s)\right)^{-1}$ and the definitions of $\boldsymbol{G}_{\operatorname{p}}$ and $\boldsymbol{G}_{\operatorname{v}}$ in eq. 18. Likewise, applying the second resolvent identity eq. 19b reveals that
$\mathbb{K}_{\boldsymbol{k},\boldsymbol{j}}$ $=\boldsymbol{I}_{k,p}^{\mkern-1.5mu\mathsf{T}}\left({\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{K}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\right)\boldsymbol{I}_{j,m}$
$=\varphi_{k}\varrho_{j}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{k}\boldsymbol{C}_{\operatorname{v}}\right)\boldsymbol{\varphi}(\mathfrak{i}\vartheta_{k})^{-1}\boldsymbol{K}\boldsymbol{\varphi}(\mathfrak{i}\zeta_{j})^{-1}\boldsymbol{B}_{\operatorname{u}}$
$=\varphi_{k}\varrho_{j}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{k}\boldsymbol{C}_{\operatorname{v}}\right)\frac{\boldsymbol{\phi}\left(h(\mathfrak{i}\vartheta_{k})\right)^{-1}\boldsymbol{K}\boldsymbol{\phi}\left(h(\mathfrak{i}\zeta_{j})\right)^{-1}}{d(\mathfrak{i}\vartheta_{k})d(\mathfrak{i}\zeta_{j})}\boldsymbol{B}_{\operatorname{u}}$
$=\varphi_{k}\varrho_{j}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{k}\boldsymbol{C}_{\operatorname{v}}\right)\frac{h(\mathfrak{i}\vartheta_{k})\boldsymbol{\phi}\left(h(\mathfrak{i}\vartheta_{k})\right)^{-1}-h(\mathfrak{i}\zeta_{j})\boldsymbol{\phi}\left(h(\mathfrak{i}\zeta_{j})\right)^{-1}}{d(\mathfrak{i}\vartheta_{k})d(\mathfrak{i}\zeta_{j})\left(h(\mathfrak{i}\vartheta_{k})-h(\mathfrak{i}\zeta_{j})\right)}\boldsymbol{B}_{\operatorname{u}}$
$=\frac{\varphi_{k}\varrho_{j}}{d(\mathfrak{i}\vartheta_{k})d(\mathfrak{i}\zeta_{j})}\frac{n(\mathfrak{i}\vartheta_{k})\boldsymbol{G}(\mathfrak{i}\vartheta_{k})-n(\mathfrak{i}\zeta_{j})\left(\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j})+\frac{\vartheta_{k}}{\zeta_{j}}\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j})\right)}{h(\mathfrak{i}\vartheta_{k})-h(\mathfrak{i}\zeta_{j})},$
thus proving the final remaining formula eq. 22b. ∎
In the formulae eqs. 22a and 22b, the appearing term $\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j})+(\vartheta_{k}/\zeta_{j})\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j})$ looks contrived at first glance. However, this is an artifact of Theorem 1 being stated as generally as possible to include position and velocity outputs simultaneously; neither is necessarily required. In realistic applications, it is usually the case that either $\boldsymbol{C}_{\operatorname{p}}$ or $\boldsymbol{C}_{\operatorname{v}}$ is identically zero. Then, the term will resolve to
$\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j})+\frac{\vartheta_{k}}{\zeta_{j}}\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j})=\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j})=\boldsymbol{G}(\mathfrak{i}\zeta_{j})\quad\text{if}\quad\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{0},$
which is exactly the second-order transfer function eq. 2, or to
$\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j})+\frac{\vartheta_{k}}{\zeta_{j}}\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j})=\frac{\vartheta_{k}}{\zeta_{j}}\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j})=\frac{\vartheta_{k}}{\zeta_{j}}\boldsymbol{G}(\mathfrak{i}\zeta_{j})\quad\text{if}\quad\boldsymbol{C}_{\operatorname{p}}=\boldsymbol{0},$
which is a re-scaling of the second-order transfer function eq. 2.
Replacing the intrusive quantities eq. 14 in Algorithm 1 with the data-based approximations eq. 21 results in a data-driven reformulation of $\mathsf{sopvBT}$ summarized in Algorithm 2, which we call second-order quadrature-based position-velocity balanced truncation ($\mathsf{soQuadpvBT}$). Outside of empirical knowledge used to design the damping coefficient functions $f(s)$ and $g(s)$, the method is entirely non-intrusive. As in [21], the quadrature-based approximations to the Gramians are never explicitly formed. Instead, they are only invoked implicitly to derive the data-based approximations eq. 21. While Algorithm 2 is formulated generally to include position and velocity outputs, as is the case with Theorem 1, both are not necessarily required; depending on the application of interest, only one of them is required. Lastly, Theorem 1 assumes that the sets of quadrature nodes used to implicitly approximate $\boldsymbol{P}_{\operatorname{p}}$ and $\boldsymbol{Q}_{\operatorname{v}}$ are distinct. In Theorem 2, we derive a special case of the formulae eq. 22 when the quadrature nodes come in complex conjugate pairs, i.e., $\mathfrak{i}\zeta_{j}=\mathfrak{i}s_{j}$ and $\mathfrak{i}\vartheta_{j}=-\mathfrak{i}s_{j}$ for all $j=1,\ldots,J$, and may overlap.
1
Input: System transfer function $\boldsymbol{G}$, position and velocity mappings $\boldsymbol{G}_{\operatorname{p}}$ and $\boldsymbol{G}_{\operatorname{v}}$, damping coefficients $f$ and $g$, left and right quadrature nodes and weights $\{(\mathfrak{i}\vartheta_{k},\varphi_{k})\}_{k=1}^{K}$ and $\{(\mathfrak{i}\zeta_{j},\varrho_{j})\}_{j=1}^{J}$, reduction order $r$.
Output: Reduced-order system matrices ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{D}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}},{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}},{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ in eq. 3.
2
3Evaluate the mappings at the quadrature nodes to obtain the data
$\left\{\Big{(}\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j}),\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j}),f(\mathfrak{i}\zeta_{j}),g(\mathfrak{i}\zeta_{j})\Big{)}\right\}_{j=1}^{J}\quad\text{and}\quad\left\{\Big{(}\boldsymbol{G}(\vartheta_{k}),f(\mathfrak{i}\vartheta_{k}),g(\mathfrak{i}\vartheta_{k})\Big{)}\right\}_{k=1}^{K},$
and construct the data matrices $\mathbb{M},\mathbb{K},\mathbb{B}_{\operatorname{u}},\mathbb{C}_{\operatorname{p}},\mathbb{C}_{\operatorname{v}}$ according to Theorem 1.
4
5Compute the singular value decomposition
$\mathbb{M}=\begin{bmatrix}{\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}_{1}&{\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}_{2}\end{bmatrix}\begin{bmatrix}\check{\boldsymbol{\Sigma}}_{1}&\boldsymbol{0}\\ \boldsymbol{0}&\check{\boldsymbol{\Sigma}}_{2}\end{bmatrix}\begin{bmatrix}{\check{\boldsymbol{Y}\mkern 2.0mu}\mkern-2.0mu}{}_{1}^{\mathsf{H}}\\[2.0pt] {\check{\boldsymbol{Y}\mkern 2.0mu}\mkern-2.0mu}{}_{2}^{\mathsf{H}}\end{bmatrix},$
for $\check{\boldsymbol{\Sigma}}_{1}\in\mathbb{R}^{r\times r}$ diagonal containing the $r$ largest singular values, $\check{\boldsymbol{\Sigma}}_{2}\in\mathbb{R}^{(pK-r)\times(mJ-r)}$ diagonal and ${\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}_{1},{\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}_{2},{\check{\boldsymbol{Y}\mkern 2.0mu}\mkern-2.0mu}{}_{1},{\check{\boldsymbol{Y}\mkern 2.0mu}\mkern-2.0mu}{}_{2}$ partitioned accordingly.
6
7Compute the reduced-order model matrices according to
${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}$ $=\boldsymbol{I}_{r},$
${\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}$ $=\check{\boldsymbol{\Sigma}}_{1}^{-1/2}{\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}_{1}^{\mathsf{H}}\mathbb{K}{\check{\boldsymbol{Y}\mkern 2.0mu}\mkern-2.0mu}{}_{1}\check{\boldsymbol{\Sigma}}_{1}^{-1/2},$
[…]
${\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}$ $=\check{\boldsymbol{\Sigma}}_{1}^{-1/2}{\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}_{1}^{\mathsf{H}}\mathbb{B}_{\operatorname{u}},$
${\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}}$ $=\mathbb{C}_{\operatorname{p}}{\check{\boldsymbol{Y}\mkern 2.0mu}\mkern-2.0mu}{}_{1}\check{\boldsymbol{\Sigma}}_{1}^{-1/2},$
${\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ $=\mathbb{C}_{\operatorname{v}}{\check{\boldsymbol{Y}\mkern 2.0mu}\mkern-2.0mu}{}_{1}\check{\boldsymbol{\Sigma}}_{1}^{-1/2}.$
Algorithm 2 Quadrature-based $\mathsf{sopvBT}$ ($\mathsf{soQuadpvBT}$).
Remark 3.
It is worth highlighting that the matrices in eq. 21 are highly similar to the second-order Loewner and shifted Loewner matrices that appear in the extension of the interpolatory Loewner framework for second-order systems developed in [32]. Suppose that $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{0}$ and that the system eq. 1 has Rayleigh damping eq. 6. Given two disjoint sets of interpolation points $\lambda_{1},\ldots,\lambda_{K}\in\mathbb{C}$ and $\mu_{1},\ldots,\mu_{J}\in\mathbb{C}$, the second-order Loewner and shifted Loewner matrices $\mathbb{L}\in\mathbb{C}^{pK\times mJ}$ and $\mathbb{L}^{\operatorname{s}}\in\mathbb{C}^{pK\times mJ}$ are defined block-elementwise as
$\mathbb{L}_{\boldsymbol{k},\boldsymbol{j}}=\frac{d(\lambda_{k})\boldsymbol{G}(\lambda_{k})-d(\mu_{j})\boldsymbol{G}(\mu_{j})}{h(\lambda_{k})-h(\mu_{j})},\quad\mathbb{L}^{\operatorname{s}}_{\boldsymbol{k},\boldsymbol{j}}=\frac{n(\lambda_{k})\boldsymbol{G}(\lambda_{k})-n(\mu_{j})\boldsymbol{G}(\mu_{j})}{h(\lambda_{k})-h(\mu_{j})},$ (24)
where $n$, $d$ and $h$ are defined as in eq. 20 for $f(s)=\alpha$ and $g(s)=\beta$; see [32, Def. 2]. If $\boldsymbol{C}_{\operatorname{p}}=\boldsymbol{0}$, and we take $\mu_{j}=\mathfrak{i}\zeta_{j}$ and $\lambda_{k}=\mathfrak{i}\vartheta_{k}$, then the matrices eq. 21, which appear in $\mathsf{soQuadpvBT}$, are related to the second-order Loewner and shifted Loewner matrices eq. 24 by diagonal scalings:
$\mathbb{M}$ $=-\operatorname{diag}\left(\varphi_{1}d(\mathfrak{i}\vartheta_{1})^{-1},\ldots,\varphi_{K}d(\mathfrak{i}\vartheta_{K})^{-1}\right)\mathbb{L}\operatorname{diag}\left(\varrho_{1}d(\mathfrak{i}\zeta_{1})^{-1},\ldots,\varrho_{J}d(\mathfrak{i}\zeta_{J})^{-1}\right),$
$\mathbb{K}$ $=\hphantom{-}\operatorname{diag}\left(\varphi_{1}d(\mathfrak{i}\vartheta_{1})^{-1},\ldots,\varphi_{K}d(\mathfrak{i}\vartheta_{K})^{-1}\right)\mathbb{L}^{\operatorname{s}}\operatorname{diag}\left(\varrho_{1}d(\mathfrak{i}\zeta_{1})^{-1},\ldots,\varrho_{J}d(\mathfrak{i}\zeta_{J})^{-1}\right).$
In [32], the matrices eq. 24 are used to construct rational interpolants of second-order systems from data; see [32, Thm. 3] for further details. The input and output matrices of this rational interpolant are those in eq. 22c multiplied from the left and right, respectively, by the diagonal scalings given above. We also highlight that the derivation of the diagonally scaled Loewner matrices in the proof of Theorem 1 is different from that of [32, Thm. 3], which uses a parameterization of interpolating systems eq. 3 based on matrix equations. The tools and techniques used in the proof of Theorem 1 can be applied to derive this generalization of the Loewner framework from [32], as well. Significantly, this extends the applicability of the framework to second-order systems eq. 1 with the generalized proportional damping model eq. 5, as well as mixed position and velocity outputs.
Remark 4.
In [44], the authors re-derive the formulas from the dissertation [36] for the case of second-order velocity balanced truncation. In contrast to [44], our work not only considers the position-velocity balanced truncation approach but is also presented for a broader class of dynamical systems, namely multiple-input/multiple-output second-order systems with position as well as velocity outputs and with the generalized proportional damping model eq. 5. Additionally, we note that most of the following considerations in Sections 4 and 5 have not been discussed for the velocity balancing approach in [44].
4 Theoretical considerations
In this section, we consider theoretical implications of the $\mathsf{soQuadpvBT}$ framework including the preservation of asymptotic stability and the error between the data-driven terms and their intrusive counterparts.
4.1 Stability preservation
For the present discussion, we allow the matrices in eq. 1 to be complex valued. In the intrusive setting, $\mathsf{sopvBT}$ preserves asymptotic stability when the underlying full-order model eq. 1 is Hermitian, i.e.,
$\boldsymbol{M}=\boldsymbol{M}^{\mathsf{H}},\quad\boldsymbol{D}=\boldsymbol{D}^{\mathsf{H}},\quad\boldsymbol{K}=\boldsymbol{K}^{\mathsf{H}},\quad\boldsymbol{B}_{\operatorname{u}}=\boldsymbol{C}_{\operatorname{p}}^{\mathsf{H}}\quad\text{and}\quad\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{0},$ (25)
and the system matrices $\boldsymbol{M},\boldsymbol{D},\boldsymbol{K}$ are Hermitian positive definite ($HPD$). Obviously, a special case of eq. 25 is the system eq. 1 being state-space symmetric when the corresponding matrices are real valued. Interestingly, $\mathsf{soQuadpvBT}$ will preserve asymptotic stability if the data are generated by a Rayleigh-damped system that is Hermitian with $HPD$ mass, damping and stiffness matrices, and if the left and right quadrature nodes—those used to implicitly approximate $\boldsymbol{Q}_{\operatorname{v}}$ and $\boldsymbol{P}_{\operatorname{p}}$, respectively—are complex conjugate pairs of each other, i.e., $\mathfrak{i}\zeta_{i}=\mathfrak{i}s_{i}$ and $\mathfrak{i}\vartheta_{i}=-\mathfrak{i}s_{i}$. If the $s_{i}$’s are not all strictly positive or negative, then at least some of the nodes will be equivalent, i.e., $\zeta_{j}=s_{j}=-s_{k}=\vartheta_{k}$ for some $j,k$, which is not allowed under Theorem 1. To prove the stability result, we first derive formulae for the data matrices eq. 21 in this special case when the quadrature nodes underlying the approximations to $\boldsymbol{Q}_{\operatorname{v}}$ and $\boldsymbol{P}_{\operatorname{p}}$ come in complex conjugate pairs and may overlap, with $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{0}$.
Theorem 2.
Let the functions $d\colon\mathbb{C}\to\mathbb{C}$, $n\colon\mathbb{C}\to\mathbb{C}$ and $h\colon\mathbb{C}\to\mathbb{C}$ be defined as in eq. 20. Suppose the left and right quadrature rules in eqs. 16 and 17 are complex conjugate pairs, i.e., $\zeta_{i}=s_{i}$, $\vartheta_{i}=-s_{i}$ and $\eta_{i}=\varrho_{i}=\vartheta_{i}$ for $i=1,\ldots,N$ and $N=J=K$, and the quadrature rules are such that $d(\mathfrak{i}s_{i})\neq 0$ and $\boldsymbol{\varphi}(\mathfrak{i}s_{j})$ is nonsingular for all $i=1,\ldots,N$. Let the quadrature-based factors ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}$ and ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ be given as in eqs. 16 and 17 and let the matrices $\mathbb{M},\mathbb{K}\in\mathbb{C}^{pK\times mJ}$, $\mathbb{B}_{\operatorname{u}}\in\mathbb{C}^{pK\times m}$ and $\mathbb{C}_{\operatorname{p}}\in\mathbb{C}^{p\times mJ}$ be as in eq. 21. Then, the matrices $\mathbb{M}$, $\mathbb{K}$, $\mathbb{B}_{\operatorname{u}}$ and $\mathbb{C}_{\operatorname{p}}$ are given block-entrywise by
$\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}$ =$=\begin{cases}\displaystyle-\frac{\eta_{k}\eta_{j}}{d(\mathfrak{i}s_{j})^{2}}\frac{\operatorname{d}}{\operatorname{d}s}\Big{(}d(s)\boldsymbol{G}(s)\Big{)}\bigg{|}_{s=\mathfrak{i}s_{j}}\left(\frac{\operatorname{d}}{\operatorname{d}s}h(s)\bigg{|}_{s=\mathfrak{i}s_{j}}\right)^{-1}&\text{if}\quad\mathfrak{i}s_{j}=-\mathfrak{i}s_{k},\\[14.22636pt] \displaystyle-\frac{\eta_{k}\eta_{j}}{d(-\mathfrak{i}s_{k})d(\mathfrak{i}s_{j})}\frac{d(-\mathfrak{i}s_{k})\boldsymbol{G}(-\mathfrak{i}s_{k})-d(\mathfrak{i}s_{j})\boldsymbol{G}(\mathfrak{i}s_{j})}{h(-\mathfrak{i}s_{k})-h(\mathfrak{i}s_{j})}&\text{if}\quad\mathfrak{i}s_{j}\neq-\mathfrak{i}s_{k},\end{cases}$ (26a)
$\mathbb{K}_{\boldsymbol{k},\boldsymbol{j}}$ =$=\begin{cases}\displaystyle\frac{\eta_{k}\eta_{j}}{d(\mathfrak{i}s_{j})^{2}}\frac{\operatorname{d}}{\operatorname{d}s}\Big{(}n(s)\boldsymbol{G}(s)\Big{)}\bigg{|}_{s=\mathfrak{i}s_{j}}\left(\frac{\operatorname{d}}{\operatorname{d}s}h(s)\bigg{|}_{s=\mathfrak{i}s_{j}}\right)^{-1}&\text{if}\quad\mathfrak{i}s_{j}=-\mathfrak{i}s_{k},\\[14.22636pt] \displaystyle\frac{\eta_{k}\eta_{j}}{d(-\mathfrak{i}s_{k})d(\mathfrak{i}s_{j})}\frac{n(-\mathfrak{i}s_{k})\boldsymbol{G}(-\mathfrak{i}s_{k})-n(\mathfrak{i}s_{j})\boldsymbol{G}(\mathfrak{i}s_{j})}{h(-\mathfrak{i}s_{k})-h(\mathfrak{i}s_{j})}&\text{if}\quad\mathfrak{i}s_{j}\neq-\mathfrak{i}s_{k},\end{cases}$ (26b)
$\big{(}\mathbb{B}_{\operatorname{u}}\big{)}_{\boldsymbol{k},:}$ $=\eta\boldsymbol{G}(-\mathfrak{i}s_{k})\quad\text{and}\quad\big{(}\mathbb{C}_{\operatorname{p}}\big{)}_{:,\boldsymbol{j}}=\eta\boldsymbol{G}(\mathfrak{i}s_{j}),$ (26c)
where $\boldsymbol{G}$ is the second-order transfer function eq. 2.
Proof.
The claims in eq. 26c, as well as eqs. 26a and 26b for the case of $\mathfrak{i}s_{j}\neq-\mathfrak{i}s_{k}$ follow directly from Theorem 1 under the stated assumptions since $\zeta_{j}=s_{j}$ and $\vartheta_{k}=-s_{k}$. To prove eq. 26a in the $\mathfrak{i}s_{j}=-\mathfrak{i}s_{k}$ case, we observe that
$\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}=\boldsymbol{I}_{k,p}^{\mkern-1.5mu\mathsf{T}}\left({\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{M}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\right)\boldsymbol{I}_{j,m}=\eta_{k}\eta_{j}\boldsymbol{C}_{\operatorname{p}}\boldsymbol{\varphi}(\mathfrak{i}s_{j})^{-1}\boldsymbol{M}\boldsymbol{\varphi}(\mathfrak{i}s_{j})^{-1}\boldsymbol{B}_{\operatorname{u}}$
holds. Let $\varepsilon>0$; using the same logic as in the proof of Theorem 1, we have that
$\boldsymbol{C}_{\operatorname{p}}\boldsymbol{\varphi}(\mathfrak{i}s_{j})^{-1}\boldsymbol{M}\boldsymbol{\varphi}(\mathfrak{i}s_{j})^{-1}\boldsymbol{B}_{\operatorname{u}}$
$=\lim\limits_{\varepsilon\to 0}\boldsymbol{C}_{\operatorname{p}}\boldsymbol{\varphi}\big{(}\mathfrak{i}(s_{j}+\varepsilon)\big{)}^{-1}\boldsymbol{M}\boldsymbol{\varphi}(\mathfrak{i}s_{j})^{-1}$
$=-\lim\limits_{\varepsilon\to 0}\frac{1}{d\big{(}\mathfrak{i}(s_{j}+\varepsilon)\big{)}d(\mathfrak{i}s_{j})}\frac{d\big{(}\mathfrak{i}(s_{j}+\varepsilon)\big{)}\boldsymbol{G}\big{(}\mathfrak{i}(s_{j}+\varepsilon)\big{)}-d(\mathfrak{i}s_{j})\boldsymbol{G}(\mathfrak{i}s_{j})}{h\big{(}\mathfrak{i}(s_{j}+\varepsilon)\big{)}-h(\mathfrak{i}s_{j})}$
$=-\frac{1}{d(\mathfrak{i}s_{j})^{2}}\lim\limits_{\varepsilon\to 0}\frac{d\big{(}\mathfrak{i}(s_{j}+\varepsilon)\big{)}\boldsymbol{G}\big{(}\mathfrak{i}(s_{j}+\varepsilon)\big{)}-d(\mathfrak{i}s_{j})\boldsymbol{G}(\mathfrak{i}s_{j})}{\mathfrak{i}(s_{j}+\varepsilon)-\mathfrak{i}s_{j}}\frac{\mathfrak{i}(s_{j}+\varepsilon)-\mathfrak{i}s_{j}}{h\big{(}\mathfrak{i}(s_{j}+\varepsilon)\big{)}-h(\mathfrak{i}s_{j})}$
$=-\frac{1}{d(\mathfrak{i}s_{j})^{2}}\frac{\operatorname{d}}{\operatorname{d}s}\Big{(}d(s)\boldsymbol{G}(s)\Big{)}\bigg{|}_{s=\mathfrak{i}s_{j}}\left(\frac{\operatorname{d}}{\operatorname{d}s}h(s)\bigg{|}_{s=\mathfrak{i}s_{j}}\right)^{-1},$
by the limit definition of the derivative. This proves eq. 26a when $\mathfrak{i}s_{j}=-\mathfrak{i}s_{k}$ by multiplication with $\eta_{k}\eta_{j}$. Using the same arguments, the other remaining formula eq. 26b can be shown for $\mathfrak{i}s_{j}=-\mathfrak{i}s_{k}$, which concludes the proof. ∎
Algorithm 2 can be used with the setup and assumptions outlined in Theorem 2 to compute reduced-order models via $\mathsf{soQuadpvBT}$. The substantial difference to using Theorem 1 is that the data matrices in eq. 21 are constructed according to the formulae in eq. 26 instead of eq. 22. In this case, the matrices in eq. 21 may require information about the derivative of $\boldsymbol{G}$ in the off-diagonal entries if the chosen quadrature nodes are such that $\mathfrak{i}s_{j}=-\mathfrak{i}s_{k}$. This will indeed be the case for the realification procedure that we describe in Section 5.1.
Suppose the assumptions of Theorem 2 are satisfied and that the transfer function data $\boldsymbol{G}(\mathfrak{i}s_{j})$ are generated by a Hermitian system eq. 25 with $HPD$ mass, damping and stiffness matrices. Because the underlying system is Hermitian, it follows that the quadrature-based factors eqs. 16 and 17 satisfy ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}={\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ since $\mathfrak{i}\zeta_{j}=\mathfrak{i}s_{j}$ and $\mathfrak{i}\vartheta_{k}=-\mathfrak{i}s_{k}$. Thus, the matrices $\mathbb{M}$ and $\mathbb{K}$ constructed according to Theorem 2 satisfy
$\mathbb{M}={\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}^{\mathsf{H}}\boldsymbol{M}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}=\mathbb{M}^{\mathsf{H}}\quad\text{and}\quad\mathbb{K}={\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}^{\mathsf{H}}\boldsymbol{K}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}=\mathbb{K}^{\mathsf{H}}$
by the property eq. 25 of the underlying system. Moreover, $\mathbb{B}_{\operatorname{u}}=\mathbb{C}_{\operatorname{p}}^{\mathsf{H}}$ holds by construction. We claim that a reduced-order model eq. 3 computed via Algorithm 2 using these Hermitian data matrices has a similar symmetry structure as the original underlying system with $HPD$ mass, damping and stiffness matrices. Obviously ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}=\boldsymbol{I}_{r}$ is $HPD$ by construction. Because $\mathbb{M}$ is Hermitian, its singular value decomposition is so that the singular vectors satisfy ${\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}={\check{\boldsymbol{Y}\mkern 2.0mu}\mkern-2.0mu}{}$. Then, for any $\boldsymbol{z}\in\mathbb{C}^{Nm}$, it holds that
$\boldsymbol{z}^{\mathsf{H}}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{z}=\boldsymbol{z}^{\mathsf{H}}\check{\boldsymbol{\Sigma}}_{1}^{-1/2}{\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}_{1}^{\mathsf{H}}\mathbb{K}{\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}_{1}\check{\boldsymbol{\Sigma}}_{1}^{-1/2}\boldsymbol{z}=\boldsymbol{w}^{\mathsf{H}}\boldsymbol{K}\boldsymbol{w}>0,$
with $\boldsymbol{w}={\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}{\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}_{1}\check{\boldsymbol{\Sigma}}_{1}^{-1/2}\boldsymbol{z}$, and the fact that $\boldsymbol{w}^{\mathsf{H}}\boldsymbol{K}\boldsymbol{w}>0$ follows from the assumption that $\boldsymbol{K}$ is $HPD$. Therefore, we have that ${\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}$ is $HPD$. Finally, if the system is Rayleigh damped, then trivially ${\widetilde{\boldsymbol{D}\mkern 2.0mu}\mkern-2.0mu}{}$ is $HPD$ because it is a nonnegative linear combination of two $HPD$ matrices and $\alpha,\beta\geq 0$ by assumption. Because Hermitian second-order systems eq. 1 with positive definite mass, damping, and stiffness matrices are asymptotically stable, this argument directly implies the following stability result.
Theorem 3.
Assume the system $\mathcal{G}$ in eq. 1 used to generate the data $\boldsymbol{G}(\mathfrak{i}s_{j})$ is Hermitian eq. 25 with positive definite mass, damping and stiffness matrices, assume that $\mathcal{G}$ has Rayleigh damping eq. 6, and assume that the data $\boldsymbol{G}(\mathfrak{i}s_{j})$ satisfy the assumptions of Theorem 2. Then, the reduced-order system computed by $\mathsf{soQuadpvBT}$ using the data matrices from Theorem 2 is asymptotically stable.
Theorem 3 may also be used conversely to check the assumptions placed on the underlying system that provides the data. If a reduced-order model eq. 3 is computed by Algorithm 2 using the data matrices eq. 26 is not asymptotically stable, then the transfer function data $\boldsymbol{G}(\mathfrak{i}s_{j})$ were not generated by a symmetric system with $HPD$ mass, damping and stiffness matrices.
4.2 Error control
The result [21, Prop. 3.2] provides error control on the data matrices that appear in $\mathsf{foQuadBT}$ in terms of their intrusive counterparts. By recognizing that the proof of [21, Prop. 3.2] makes no assumption on the underlying system structure, this result can be extended directly to our setting to derive error bounds on the matrices in Theorem 1.
Theorem 4.
Assume that the quadrature rules used to derive eqs. 16 and 17 are such that the quadrature-based approximations to $\boldsymbol{P}_{\operatorname{p}}$ and $\boldsymbol{Q}_{\operatorname{v}}$ satisfy
$\lVert\boldsymbol{P}_{\operatorname{p}}-{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}^{\mathsf{H}}\rVert_{\operatorname{F}}\leq\frac{\delta}{1+\delta}\sigma_{\min}\left(\boldsymbol{P}_{\operatorname{p}}\right)\quad\text{and}\quad\lVert\boldsymbol{Q}_{\operatorname{v}}-{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\rVert_{\operatorname{F}}\leq\frac{\delta}{1+\delta}\sigma_{\min}\left(\boldsymbol{Q}_{\operatorname{v}}\right),$ (27)
for some $\delta\in(0,1)$, where $\sigma_{\min}(\cdot)$ denotes the minimum singular value of a matrix. Then, there exist partial isometries $\boldsymbol{\Psi}_{\operatorname{v}}\in\mathbb{C}^{pK\times n}$ and $\boldsymbol{\Psi}_{\operatorname{p}}\in\mathbb{C}^{mJ\times n}$ so that
$\lVert{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{M}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}-\boldsymbol{\Psi}_{\operatorname{v}}\big{(}\boldsymbol{L}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{M}\boldsymbol{R}_{\operatorname{p}}\big{)}\boldsymbol{\Psi}_{\operatorname{p}}^{\mathsf{H}}\rVert_{\operatorname{F}}$ $\leq 2\delta\lVert\boldsymbol{L}_{\operatorname{v}}\rVert_{2}\lVert\boldsymbol{M}\rVert_{2}\lVert\boldsymbol{R}_{\operatorname{p}}\rVert_{2},$ (28a)
$\lVert{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{K}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}-\boldsymbol{\Psi}_{\operatorname{v}}\big{(}\boldsymbol{L}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{K}\boldsymbol{R}_{\operatorname{p}}\big{)}\boldsymbol{\Psi}_{\operatorname{p}}^{\mathsf{H}}\rVert_{\operatorname{F}}$ $\leq 2\delta\lVert\boldsymbol{L}_{\operatorname{v}}\rVert_{2}\lVert\boldsymbol{K}\rVert_{2}\lVert\boldsymbol{R}_{\operatorname{p}}\rVert_{2},$ (28b)
$\lVert{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{B}-\boldsymbol{\Psi}_{\operatorname{v}}\big{(}\boldsymbol{L}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{B}\big{)}\rVert_{\operatorname{F}}$ $\leq 2\delta\lVert\boldsymbol{L}_{\operatorname{v}}\rVert_{2}\lVert\boldsymbol{B}\rVert_{2},$ (28c)
$\lVert\boldsymbol{C}_{\operatorname{p}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}-\big{(}\boldsymbol{C}_{\operatorname{p}}\boldsymbol{R}_{\operatorname{p}}\big{)}\boldsymbol{\Psi}_{\operatorname{p}}^{\mathsf{H}}\rVert_{\operatorname{F}}$ $\leq 2\delta\lVert\boldsymbol{C}_{\operatorname{p}}\rVert_{2}\lVert\boldsymbol{R}_{\operatorname{p}}\rVert_{2},$ (28d)
$\lVert\boldsymbol{C}_{\operatorname{v}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}-\big{(}\boldsymbol{C}_{\operatorname{v}}\boldsymbol{R}_{\operatorname{p}}\big{)}\boldsymbol{\Psi}_{\operatorname{p}}^{\mathsf{H}}\rVert_{\operatorname{F}}$ $\leq 2\delta\lVert\boldsymbol{C}_{\operatorname{v}}\rVert_{2}\lVert\boldsymbol{R}_{\operatorname{p}}\rVert_{2}.$ (28e)
Proof.
The result follows by carefully following the proof of [21, Prop. 3.2] in the context of Theorem 1. ∎
In general, Theorem 4 can be applied to any low-rank implementation of $\mathsf{sopvBT}$ for which the approximate factors satisfy eq. 27. For our setting, this allows us to control the error in the reduced-order model matrices via the error in the corresponding quadrature rules. Thus, for a sufficiently fine resolved quadrature rule, one can recover nearly exact $\mathsf{sopvBT}$-based reduced-order models solely from data.
5 Practical considerations
In this section, we discuss practical considerations for the implementation of Algorithm 2. First, we provide formulas for the $\mathsf{soQuadpvBT}$ method to construct real-valued reduced-order models, and then we discuss how to infer suitable damping parameters for the proportional damping model from data.
5.1 Enforcing real-valued models from complex-valued data
In the context of systems used in time-domain simulations, the underlying system eq. 1 is typically presumed to be real in the sense that the inputs, state variables, and outputs evolve in a real-dimensional vector space. Thus, there exists a state-space realization of eq. 1 given by real-valued matrices $\boldsymbol{M}$, $\boldsymbol{D}$, $\boldsymbol{K}$, $\boldsymbol{B}_{\operatorname{u}}$, $\boldsymbol{C}_{\operatorname{p}}$ and $\boldsymbol{C}_{\operatorname{v}}$. When such a real system generates the transfer function data used in Algorithm 2, it is possible to modify $\mathsf{soQuadpvBT}$ to produce a real-valued realization of the reduced-order model. In the subsequent discussion, we describe how to achieve the construction of such real-valued realizations. To this end, we require that the frequency-dependent damping coefficients $f$ and $g$ in eq. 5 to satisfy
$\overline{f(s)}$ $=f(\overline{s}),\quad\overline{g(s)}=g(\overline{s})\quad\text{so that}\quad$ (29a)
$\overline{n(s)}$ $=n(\overline{s}),\quad\overline{d(s)}=d(\overline{s}),\quad\overline{h(s)}=h(\overline{s}),$ (29b)
for all $s\in\mathbb{C}$, where $n$, $d$ and $h$ are defined as in eq. 20. The assumption eq. 29 is trivially satisfied for Rayleigh damping eq. 6 because $\alpha,\beta\in\mathbb{R}$, and notably not satisfied for structural damping eq. 7 because $\overline{g(s)}=-\mathfrak{i}\tfrac{\overline{\eta}}{\overline{s}}\neq\mathfrak{i}\tfrac{\overline{\eta}}{\overline{s}}=g(\overline{s})$. Although this is expected; indeed, a structurally-damped system can always be expressed by a realization with zero damping and complex-valued stiffness $\left(1+\mathfrak{i}\eta\right)\boldsymbol{K}$.
For simplicity of presentation, we assume that both quadrature rules use an equal number of nodes, $N=J=K$, and that $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{0}$; the construction that we describe in the following also applies to the general case. Suppose that a symmetric quadrature rule is used to generate the approximate factors eqs. 16 and 17, i.e., the nodes are symmetrically distributed according to
$\vartheta_{1}$ $<\vartheta_{3}<\ldots<\vartheta_{N-1}<0<\vartheta_{2}<\vartheta_{4}<\ldots<\vartheta_{N},$ (30a)
$\zeta_{1}$ $<\zeta_{3}<\ldots<\zeta_{N-1}<0<\zeta_{2}<\zeta_{4}<\ldots<\zeta_{N},$ (30b)
and satisfy $\vartheta_{i}=-\vartheta_{i+1}$ and $\zeta_{i}=-\zeta_{i+1}$, for $i=1,3,\ldots,N-1$. We assume additionally that the corresponding quadrature weights are computed such that $\varrho_{i}=\varrho_{i+1}$ and $\varphi_{i}=\varphi_{i+1}$ for all $i=1,3,\ldots,N-1$. This symmetric ordering scheme ensures that the transfer function evaluations at these nodes come in complex conjugate pairs, that is, $\overline{\boldsymbol{G}(\mathfrak{i}\vartheta_{i})}=\boldsymbol{G}(\mathfrak{i}\vartheta_{i+1})$ and $\overline{\boldsymbol{G}(\mathfrak{i}\zeta_{i})}=\boldsymbol{G}(\mathfrak{i}\zeta_{i+1})$, for $i=1,3\ldots,N-1$. From this conjugate symmetry and the assumption eq. 29, the $2p\times 2m$-block entries of $\mathbb{M}$ corresponding to the nodes $(\mathfrak{i}\vartheta_{k},-\mathfrak{i}\vartheta_{k})$ and $(\mathfrak{i}\zeta_{j},-\mathfrak{i}\zeta_{j})$ satisfy
$\begin{bmatrix}\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}&\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}+\boldsymbol{1}}\\ \mathbb{M}_{\boldsymbol{k}+\boldsymbol{1},\boldsymbol{j}}&\mathbb{M}_{\boldsymbol{k}+\boldsymbol{1},\boldsymbol{j}+\boldsymbol{1}}\end{bmatrix}$
$=-\varphi_{k}\varrho_{j}\begin{bmatrix}\displaystyle\frac{d(\mathfrak{i}\vartheta_{k})\boldsymbol{G}(\mathfrak{i}\vartheta_{k})-d(\mathfrak{i}\zeta_{j})\boldsymbol{G}(\mathfrak{i}\zeta_{j})}{d(\mathfrak{i}\vartheta_{k})d(\mathfrak{i}\zeta_{j})\left(h(\mathfrak{i}\vartheta_{k})-h(\mathfrak{i}\zeta_{j})\right)}&\displaystyle\frac{d(\mathfrak{i}\vartheta_{k})\boldsymbol{G}(\mathfrak{i}\vartheta_{k})-\overline{d(\mathfrak{i}\zeta_{j})\boldsymbol{G}(\mathfrak{i}\zeta_{j})}}{d(\mathfrak{i}\vartheta_{k})\overline{d(\mathfrak{i}\zeta_{j})}\left(h(\mathfrak{i}\vartheta_{k})-\overline{h(\mathfrak{i}\zeta_{j})}\right)}\\[3.0pt] \displaystyle\frac{\overline{d(\mathfrak{i}\vartheta_{k})\boldsymbol{G}(\mathfrak{i}\vartheta_{k})}-d(\mathfrak{i}\zeta_{j})\boldsymbol{G}(\mathfrak{i}\zeta_{j})}{\overline{d(\mathfrak{i}\vartheta_{k})}d(\mathfrak{i}\zeta_{j})\left(\overline{h(\mathfrak{i}\vartheta_{k})}-h(\mathfrak{i}\zeta_{j})\right)}&\displaystyle\frac{\overline{d(\mathfrak{i}\vartheta_{k})}\overline{\boldsymbol{G}(\mathfrak{i}\vartheta_{k})}-\overline{d(\mathfrak{i}\zeta_{j})}\overline{\boldsymbol{G}(\mathfrak{i}\zeta_{j})}}{\overline{d(\mathfrak{i}\vartheta_{k})}\overline{d(\mathfrak{i}\zeta_{j})}\left(\overline{h(\mathfrak{i}\vartheta_{k})}-\overline{h(\mathfrak{i}\zeta_{j})}\right)}\end{bmatrix},$
where $\boldsymbol{k}$ and $\boldsymbol{k}+\boldsymbol{1}$ refer to the indices $(k-1)p+1:kp$ and $kp+1:(k+1)p$ and similarly for $\boldsymbol{j}$ and $\boldsymbol{j}+\boldsymbol{1}$. From the above, we observe that $\overline{\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}}=\mathbb{M}_{\boldsymbol{k}+\boldsymbol{1},\boldsymbol{j}+\boldsymbol{1}}$ and $\overline{\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}+\boldsymbol{1}}}=\mathbb{M}_{\boldsymbol{k}+\boldsymbol{1},\boldsymbol{j}}$. Define the unitary matrix $\boldsymbol{J}_{\ell}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\frac{1}{\sqrt{2}}\begin{bmatrix}\boldsymbol{I}_{\ell}&-\mathfrak{i}\boldsymbol{I}_{\ell}\\ \boldsymbol{I}_{\ell}&\hphantom{-}\mathfrak{i}\boldsymbol{I}_{\ell}\end{bmatrix}\in\mathbb{C}^{2\ell\times 2\ell}$ for some positive integer $\ell\in\mathbb{N}$. One can readily verify that
$\boldsymbol{J}_{p}^{\mathsf{H}}\begin{bmatrix}\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}&\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}+\boldsymbol{1}}\\ \mathbb{M}_{\boldsymbol{k}+\boldsymbol{1},\boldsymbol{j}}&\mathbb{M}_{\boldsymbol{k}+\boldsymbol{1},\boldsymbol{j}+\boldsymbol{1}}\end{bmatrix}\boldsymbol{J}_{m}=\begin{bmatrix}\hphantom{-}\operatorname{Re}(\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}+\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}+\boldsymbol{1}})&\operatorname{Im}(\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}-\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}+\boldsymbol{1}})\\ -\operatorname{Im}(\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}+\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}+\boldsymbol{1}})&\operatorname{Re}(\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}-\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}+\boldsymbol{1}})\end{bmatrix},$
which is real valued. Thus, the matrix $\mathbb{M}$ in eq. 22a is unitarily equivalent to a real-valued matrix $\widetilde{\mathbb{M}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}(\boldsymbol{I}_{N/2}\otimes\boldsymbol{J}_{p}^{\mathsf{H}})\mathbb{M}(\boldsymbol{I}_{N/2}\otimes\boldsymbol{J}_{m})\in\mathbb{R}^{pK\times mJ}$, where $\otimes$ denotes the matrix Kronecker product [20, Chap. 12.3]. Similarly, the matrices $\mathbb{K}$, $\mathbb{B}_{\operatorname{u}}$ and $\mathbb{C}_{\operatorname{p}}$ in eqs. 22b and 22c are unitarily equivalent to real-valued matrices as well. Namely, one can verify that
$\widetilde{\mathbb{K}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}(\boldsymbol{I}_{N/2}\otimes\boldsymbol{J}_{p}^{\mathsf{H}})\mathbb{K}(\boldsymbol{I}_{N/2}\otimes\boldsymbol{J}_{m}),\quad\widetilde{\mathbb{B}}_{\operatorname{u}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}(\boldsymbol{I}_{N/2}\otimes\boldsymbol{J}_{p}^{\mathsf{H}}\big{)}\mathbb{B}_{\operatorname{u}}\quad\mbox{and}\quad\widetilde{\mathbb{C}}_{\operatorname{p}}=\mathbb{C}_{\operatorname{p}}(\boldsymbol{I}_{N/2}\otimes\boldsymbol{J}_{m})$
are all real valued by construction. Therefore, we may use these unitary transformations to obtain real-valued data matrices in Algorithm 2 and, consequently, a real-valued realizations of the reduced-order models. Lastly, we mention that this realification procedure can also be applied to the second-order Hermite Loewner matrices derived in Theorem 2. If the underlying system is real valued and the data satisfy the hypotheses of Theorem 3, then this allows one to obtain a real-valued realization of an asymptotically stable system. In this setting, it will be the case that $\zeta_{i+1}=s_{i+1}=-s_{i}=\vartheta_{i}$ and $\zeta_{i}=s_{i}=-s_{i+1}=\vartheta_{i+1}$ due to eq. 30, and so evaluations of the derivative of $\boldsymbol{G}$ will appear in the sub- and super-diagonals of $\mathbb{M}$ and $\mathbb{K}$.
5.2 Optimal choice of damping parameters
Thus far, we have assumed that the coefficient functions $f(s)$ and $g(s)$ in the generalized proportional damping model eq. 5 are given and can be freely evaluated. However, in a truly data-driven regime, these would be unknown. We discuss how these coefficients can be inferred from data in the simplified cases of Rayleigh damping eq. 6 and structural damping eq. 7; this corresponds to inferring the parameters $\alpha,\beta\in\mathbb{R}$ or $\eta\in\mathbb{R}$, respectively. In either case, we assume the data $\boldsymbol{g}_{k}=\boldsymbol{G}(s_{k})$ to be prescribed at a fixed number of points $s_{k}\in\mathbb{C}$, for $k=1,\ldots,N$; we do not assume this data to have any particular structure other than that it is generated by the true underlying system’s transfer function $\boldsymbol{G}$. The reduced-order matrices are computed according to Algorithm 2, and the goal is then to optimize the damping parameters in the $\mathsf{soQuadpvBT}$ reduced model to best fit the given data. A similar idea has been employed in the previously mentioned second-order Loewner framework [32], as well as in the data-driven modeling of time-delay systems [38]. In contrast to these works, we explicitly derive gradients of the relevant cost function with respect to the damping parameters.
To streamline the subsequent presentation, we introduce the following notation
$\boldsymbol{\varphi}_{\operatorname{R}}$ $\colon\mathbb{C}\to\mathbb{C}^{n\times n}$ with $\boldsymbol{\varphi}_{\operatorname{R}}(s,\alpha,\beta)$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}(s^{2}+s\alpha){\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}+(1+s\beta){\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{},$
$\boldsymbol{\varphi}_{\operatorname{S}}$ $\colon\mathbb{C}\to\mathbb{C}^{n\times n}$ with $\boldsymbol{\varphi}_{\operatorname{S}}(s,\eta)$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}s^{2}{\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}+(1+\mathfrak{i}\eta){\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{},$
$\mathcal{C}$ $\colon\mathbb{C}\to\mathbb{C}^{p\times n}$ with $\mathcal{C}(s)$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}}+s{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}.$
Even though $\mathsf{soQuadpvBT}$ constructs reduced models with ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}=\boldsymbol{I}_{r}$, this is not enforced in the subsequent results to make the presentation as general as possible. Thus, the minimization procedure we describe can be applied to any data-driven reduced model eq. 3 with unknown damping parameters; e.g., it can be applied to those computed by the second-order Loewner framework [32]. For the Rayleigh damped case, the optimization problem is formulated as
$\begin{split}(\alpha_{\star},\beta_{\star})&=\operatorname*{arg\,min}\limits_{(\alpha,\beta)\in\mathbb{R}^{2}}\sum\limits_{k=1}^{N}\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta),\quad\text{with}\\ \mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)&\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\left\lVert\boldsymbol{g}_{k}-\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\right\rVert_{\operatorname{F}}^{2},\end{split}$ (31)
and for the structurally damped case, we have
$\eta_{\star}=\operatorname*{arg\,min}\limits_{\eta\in\mathbb{R}}\sum\limits_{k=1}^{N}\mathcal{J}_{k,\operatorname{S}}(\eta),\quad\text{with}\quad\mathcal{J}_{k,\operatorname{S}}(\eta)\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\left\lVert\boldsymbol{g}_{k}-\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{S}}(s_{k},\eta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\right\rVert_{\operatorname{F}}^{2}.$ (32)
The gradients of the objective functions eqs. 31 and 32 are presented below.
Theorem 5.
Consider a reduced-order model of the form eq. 3. For the minimization problem eq. 31, the gradients with respect to the optimization parameters are
[…]
$\quad{}-{}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big{)}^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}\bigg{)}\bigg{)},$
[…]
$\quad{}-{}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}\boldsymbol{B}_{\operatorname{u}}\Big{)}^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}\bigg{)}\bigg{)}.$
For the minimization problem eq. 32, the gradient is given by
$\frac{\partial}{\partial{\eta}}\mathcal{J}_{k,\operatorname{S}}(\eta)$ $=2\operatorname{Re}\bigg{(}\operatorname{tr}\bigg{(}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{S}}(s_{k},\eta)^{-1}$ (34)
×$\quad{}\times{}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big{(}\boldsymbol{g}_{k}-\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{S}}(s_{k},\eta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big{)}^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{S}}(s_{k},\eta)^{-1}\bigg{)}\bigg{)}.$
Proof.
It suffices to show eq. 33b since the proofs of eqs. 33a and 34 follow analogously. Define $\mathcal{E}_{k}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{g}_{k}-\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\in\mathbb{C}^{p\times m}$ so that $\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)=\lVert\mathcal{E}_{k}\rVert_{\operatorname{F}}^{2}=\operatorname{tr}(\mathcal{E}_{k}^{\mathsf{H}}\mathcal{E}_{k})$. Now, we consider an arbitrary perturbation $\Delta_{\beta}\in\mathbb{R}$, and expand the helper function $\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta+\Delta_{\beta})$ as
[…]
$=\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)+s_{k}\Delta_{\beta}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}.$
By choosing $\Delta_{\beta}$ to be small enough such that $\lVert\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha_{k},\beta_{k})s_{k}\Delta_{\beta}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\rVert_{2}<1$, we may expand $\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta+\Delta_{\beta})^{-1}$ as a Neumann series into
$\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta+\Delta_{\beta})^{-1}$ $=\Big{(}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)-(-s_{k}\Delta_{\beta}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{})\Big{)}^{-1}$
$=\sum\limits_{j=0}^{\infty}\Big{(}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}(-s_{k}\Delta_{\beta}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{})\Big{)}^{j}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}$
$=\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}-s_{k}\Delta_{\beta}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}+\mathcal{O}(\lVert\Delta_{\beta}\rVert_{\operatorname{F}}^{2}).$
Using this expression, the objective function can be written as
$\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta+\Delta_{\beta})$
$=\lVert\mathcal{E}_{k}+s_{k}\Delta_{\beta}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\rVert_{\operatorname{F}}^{2}+\mathcal{O}(\lVert\Delta_{\beta}\rVert_{\operatorname{F}}^{2})$
$=\operatorname{tr}(\mathcal{E}_{k}^{\mathsf{H}}\mathcal{E}_{k})+2\operatorname{Re}\Big{(}s_{k}\operatorname{tr}\Big{(}\mathcal{E}_{k}^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big{)}\Big{)}\Delta_{\beta}+\mathcal{O}(\lVert\Delta_{\beta}\rVert_{\operatorname{F}}^{2})$
$=\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)+2\operatorname{Re}\Big{(}s_{k}\operatorname{tr}\Big{(}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\mathcal{E}_{k}^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}\Big{)}\Big{)}\Delta_{\beta}+\mathcal{O}(\lVert\Delta_{\beta}\rVert_{\operatorname{F}}^{2}),$
which proves eq. 33b. ∎
In a setting where the damping coefficients are unknown, the gradients in Theorem 5 can be used in any off-the-shelf optimization algorithm to minimize the cost function eq. 31 over possible values of $\alpha$ and $\beta$ simultaneously if the system is Rayleigh damped, or in the case of structural damping, to minimize eq. 32 over $\eta$. This minimization routine adds one additional step at the end of Algorithm 2. We emphasize that the reduced-order model matrices do not need to be recomputed after the optimization step; the computed parameters are only used to form the reduced-order damping term from the already computed mass and stiffness matrices.
6 Numerical Results
We demonstrate the proposed $\mathsf{soQuadpvBT}$ method on three benchmark examples of second-order dynamical systems from the literature. The reported numerical experiments have been performed on a MacBook Air with 8 GB of RAM and an Apple M2 processor running macOS Ventura version 13.4 with MATLAB 23.2.0.2515942 (R2023b) Update 7. The source codes, data, and results of the numerical experiments are available at [35].
6.1 Experimental setup
In comparison to our proposed $\mathsf{soQuadpvBT}$ approach, we consider the following methods in the numerical experiments:
$\mathsf{soLoewner}$
is the interpolatory Loewner framework for second-order systems with Rayleigh damping from [32] and highlighted in Remark 3;
$\mathsf{foQuadBT}$
is the data-driven balanced truncation for first-order systems from [21];
$\mathsf{sopvBT}$
is the intrusive position-velocity balanced truncation method for second-order systems from [34] and discussed in Section 2.3;
$\mathsf{BT}$
is the intrusive unstructured balanced truncation for first-order systems eq. 8 discussed in Section 2.2.
The intrusive $\mathsf{sopvBT}$ is included as a point of comparison for the data-driven aspect of $\mathsf{soQuadpvBT}$. The intrusive and data-driven first-order methods $\mathsf{BT}$ and $\mathsf{foQuadBT}$ are included for the comparison between our proposed structured approach and unstructured models. The software package MORLAB version 6.0 [12, 10] is used for computing the intrusive reduced models by $\mathsf{sopvBT}$ and $\mathsf{BT}$.
For the data-based approaches, we use an exponential trapezoidal quadrature rule to implicitly approximate the relevant Gramians in eqs. 12 and 9. While $\mathsf{soLoewner}$ does not typically use interpolation points derived from quadrature rules, we fix the sampling points across all the data-based approaches for consistency. For simplicity, an equal number of left and right quadrature nodes is used so that $J=K=N$, and we assume that the nodes are symmetrically distributed according to eq. 30. In each example, the quadrature rules will be applied in chosen intervals along the imaginary axis.
For the presentation of the numerical results, we use the following performance measures. For visual comparisons, we plot the magnitude of the second-order transfer function eq. 1 at discrete points on the positive imaginary axis. We also show the pointwise relative approximation errors of the transfer functions
$\operatorname{relerr}(\mathfrak{i}\omega_{k})=\frac{\lVert\boldsymbol{G}(\mathfrak{i}\omega_{k})-{\widetilde{\boldsymbol{G}}\mkern 2.0mu\mkern-2.0mu}{}(\mathfrak{i}\omega_{k})\rVert_{2}}{\lVert\boldsymbol{G}(\mathfrak{i}\omega_{k})\rVert_{2}},$ (35)
at discrete frequencies $\omega_{k}\in\Omega$ from the finite interval $\Omega=[\omega_{\operatorname{min}},\omega_{\operatorname{max}}]\subset[0,\infty)$. The specific choice of $\Omega$ varies between the examples. Additionally, we score the performance of the different methods by using local approximations to the relative $\mathcal{H}_{\infty}$ error and relative $\mathcal{H}_{2}$ error via
$\operatorname{relerr}_{\mathcal{H}_{\infty}}$ $=\frac{\max_{\omega_{k}\in\Omega}\lVert\boldsymbol{G}(\mathfrak{i}\omega_{k})-{\widetilde{\boldsymbol{G}}\mkern 2.0mu\mkern-2.0mu}{}(\mathfrak{i}\omega_{k})\rVert_{2}}{\max_{\omega_{k}\in\Omega}\lVert\boldsymbol{G}(\mathfrak{i}\omega_{k})\rVert_{2}},\quad\text{and}$ (36a)
$\operatorname{relerr}_{\mathcal{H}_{2}}$ $=\sqrt{\frac{\sum\limits_{\omega_{k}\in\Omega}\lVert\boldsymbol{G}(\mathfrak{i}\omega_{k})-{\widetilde{\boldsymbol{G}}\mkern 2.0mu\mkern-2.0mu}{}(\mathfrak{i}\omega_{k})\rVert_{\operatorname{F}}^{2}}{\sum\limits_{\omega_{k}\in\Omega}\lVert\boldsymbol{G}(\mathfrak{i}\omega_{k})\rVert_{\operatorname{F}}^{2}}}.$ (36b)
6.2 Butterfly Gyroscope
[…]
6.2.1 Known damping coefficients
For the first set of experiments with the butterfly gyroscope, we assume that the true damping coefficients $\alpha=0$ and $\beta=10^{-6}$ are known. Later on, in Section 6.2.2, we investigate the performance of $\mathsf{soQuadpvBT}$ reduced models with estimated damping coefficients $\alpha_{\star}$ and $\beta_{\star}$ computed using the optimization procedure from Section 5.2.
Figure 1: Frequency response and pointwise relative errors for $r=10$ reduced-order models of the butterfly gyroscope benchmark with known damping parameters. The data-driven balancing methods and their intrusive counterparts are indistinguishable.
Table 1: Relative $\mathcal{H}_{\infty}$ errors eq. 36a and $\mathcal{H}_{2}$ errors eq. 36b for the $r=10$ reduced-order models of the butterfly gyroscope with known damping coefficients. The smallest error is highlighted in boldface.
$\mathsf{soQuadpvBT}$ $\mathsf{soLoewner}$ $\mathsf{foQuadBT}$ $\mathsf{sopvBT}$ $\mathsf{BT}$
[…]
Surrogate models of order $r=10$ are computed using $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$, $\mathsf{foQuadBT}$, $\mathsf{sopvBT}$ and $\mathsf{BT}$. For the data-based approaches, we use $N=200$ left and right quadrature nodes. These are chosen to be interwoven logarithmically spaced points in the interval $-\mathfrak{i}[10^{4},10^{6}]\cup\mathfrak{i}[10^{4},10^{6}]$ and symmetrically distributed according to eq. 30. We use the realification process described in Section 5.1 for the $\mathsf{soQuadpvBT}$ and $\mathsf{soLoewner}$ reduced models; for $\mathsf{foQuadBT}$, a similar procedure from [21, Sec. 4.1] is applied. The frequency response and pointwise relative errors eq. 35 of the reduced-order models are presented in Figure 1. We see that the data-driven approaches $\mathsf{soQuadpvBT}$ and $\mathsf{foQuadBT}$ are visibly indistinguishable from their intrusive counterparts $\mathsf{soBT}$ and $\mathsf{BT}$. Also, we observe that each of the approaches that preserve the underlying second-order structure, namely $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$ and $\mathsf{sopvBT}$, provide very satisfactory approximations and are able to capture the response peaks of the full-order transfer function. By comparison, the unstructured methods $\mathsf{foQuadBT}$ and $\mathsf{BT}$ perform roughly two orders of magnitude worse throughout the entire frequency range, and entirely miss the response between $10^{5}$ and $10^{6}$ rad/s. This results from the fact that the first-order approximations having the same dimension as the second-order approximations cannot provide the same level of accuracy, regardless of whether that approximation is data-driven or intrusive. Table 1 shows the relative error measures eqs. 36a and 36b for the different methods. The structure-preserving methods clearly outperform the unstructured ones, and the $\mathsf{soQuadpvBT}$ and $\mathsf{soBT}$ reduced models produce very similar error values.
6.2.2 Estimating damping coefficients
Next, we employ the minimization routine from Section 5.2 to infer the Rayleigh damping coefficients from data, and use these estimated coefficients $\alpha_{\star}$ and $\beta_{\star}$ in the construction of the reduced-order damping matrix. For this set of experiments, we only consider $\mathsf{soQuadpvBT}$ to highlight the effects of the estimated coefficients on the reduced model error. The data $\boldsymbol{g}_{k}$ used to infer the damping coefficients are the evaluations of the transfer function $\boldsymbol{G}(\mathfrak{i}\vartheta_{k})$ and $\boldsymbol{G}(\mathfrak{i}\zeta_{k})$ at the left and right quadrature points. For solving the minimization problem eq. 31, we use the built-in MATLAB optimizer fminunc provided with the gradients eqs. 33a and 33b from Theorem 5. The internal minimization algorithm is the BFGS method with step and optimality tolerances both set to $10^{-8}$.
Figure 2: Frequency response and pointwise relative errors for the $r=10$ reduced-order $\mathsf{soQuadpvBT}$ models of the butterfly gyroscope using the true and estimated damping coefficients. The reduced-order models with inferred damping coefficients produce nearly identical results compared to the model with known parameters.
Table 2: Estimated Rayleigh damping coefficients $(\alpha_{\star},\beta_{\star})$ for the butterfly gyroscope example using different initializations $(\alpha_{i},\beta_{i})$.
$(\alpha_{1},\beta_{1})=(10^{-4},10^{-4})$ $(\alpha_{2},\beta_{2})=(0,0)$ $(\alpha_{3},\beta_{3})=(10^{-1},10^{-1})$
[…]
The minimization is applied to three different pairs of initial values for the Rayleigh damping coefficients, namely $(\alpha_{1},\beta_{1})=(10^{-4},10^{-4})$, $(\alpha_{2},\beta_{2})=(0,0)$ and $(\alpha_{3},\beta_{3})=(10^{-1},10^{-1})$, and the resulting local optima are shown in Table 2. Then, we used $\mathsf{soQuadpvBT}$ to compute four different reduced-order models based on the true damping parameters $(\alpha,\beta)=(0,10^{-6})$ and the three sets of inferred coefficients from Table 2. The frequency response and pointwise relative errors of these reduced models are all shown in Figure 2. The pointwise relative errors in Figure 2 reveal that the $\mathsf{soQuadpvBT}$ reduced models that use the estimated $\alpha_{\star}$ and $\beta_{\star}$ are nearly indistinguishable from the reduced model that incorporates the true $\alpha=0$ and $\beta=10^{-6}$. This shows that we can use $\mathsf{soQuadpvBT}$ to obtain suitable second-order reduced models as well as suitable damping coefficients using only given data.
6.2.3 Hermitian approach and stability preservation
Finally, we test the construction formulas from Theorem 2 to validate the stability result in Theorem 3. To make the system symmetric, we modify the input and position-output matrices to be $\boldsymbol{B}_{\operatorname{u}}=\boldsymbol{C}_{\operatorname{p}}=\tfrac{1}{12}\boldsymbol{1}_{n}$, where $\boldsymbol{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones of length $n$. For data generation, we use $N=200$ logarithmically spaced points $s_{i}$ in the interval $-\mathfrak{i}[10^{4},10^{6}]\cup\mathfrak{i}[10^{4},10^{6}]$ and take the left and right quadrature nodes to be such that $\vartheta_{k}=-s_{k}$ and $\zeta_{j}=s_{j}$, which are symmetrically distributed according to eq. 30, and the weights satisfying $\eta_{i}=\varphi_{i}=\varrho_{i}$. Specifically, this organizational scheme implies that $s_{i}=-s_{i+1}$, for $i=1,3,\ldots,N-1$ and so the super and sub-diagonal entries of the matrices $\mathbb{M}$ and $\mathbb{K}$ require derivatives of $\boldsymbol{G}$ as in eq. 26.
Figure 3: Frequency response and pointwise relative errors for $r=20$ $\mathsf{soQuadpvBT}$ reduced-order models of the butterfly gyroscope using the data matrices from Theorem 1 and Theorem 2. The stability preserving model performs as well as the standard approach in terms of accuracy.
Table 3: Stability of reduced models computed by $\mathsf{soQuadpvBT}$ using Theorems 1 and 2 for varying orders $r=10,12,\ldots,20$. Check marks indicate asymptotic stability, and all computed reduced-order models turned out to be asymptotically stable.
$r=10$ $r=12$ $r=14$ $r=16$ $r=18$ $r=20$
$\mathsf{soQuadpvBT}$ (Theorem 1)
$\mathsf{soQuadpvBT}$ (Theorem 2)
For varying reduction orders $r=10,12,\ldots,20$, we compute reduced-order models via $\mathsf{soQuadpvBT}$ using the setup just described and the data matrices from Theorem 2, and for the comparison, we also compute reduced-order models using the setup of Section 6.2.1 and the data matrices in Theorem 1. In all cases, we employ the realification procedure from Section 5.1 to compute real-valued surrogates. The response and pointwise relative errors eq. 35 of the $r=20$ reduced-order are plotted in Figure 3. As illustrated by the figure, the $\mathsf{soQuadpvBT}$ reduced models using either set of data matrices provide very satisfactory approximations over the entire frequency range, and are nearly indistinguishable from each other. In Table 3, we report the stability properties of the reduced-order models. As expected from Theorem 3, the $\mathsf{soQuadpvBT}$ models constructed via Theorem 2 are all asymptotically stable by construction. Although not guaranteed, the $\mathsf{soQuadpvBT}$ reduced models constructed using Theorem 1 turned out to be asymptotically stable, as well.
6.3 Plate with tuned vibration absorbers
[…]
Figure 4: Frequency response and pointwise relative errors for $r=75$ reduced-order models of the plate with TVAs. The structured approaches clearly outperform the unstructured $\mathsf{foQuadBT}$ for most frequencies by about five orders of magnitude.
Table 4: Relative $\mathcal{H}_{\infty}$ errors eq. 36a and $\mathcal{H}_{2}$ errors eq. 36b for $r=75$ reduced-order models of the plate with TVAs. The smallest error is highlighted in boldface.
$\mathsf{soQuadpvBT}$ $\mathsf{soLoewner}$ $\mathsf{foQuadBT}$
[…]
In this example, we only employ the data-driven methods $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$ and $\mathsf{foQuadBT}$ due to the lack of available matrix equation solvers for the complex-valued coefficient matrices. Furthermore, we do not employ the realification process here because the underlying hysteretic damping leads to the full-order model being complex valued. For generating the data, we choose $N=250$ left and right quadrature nodes that are interwoven linearly spaced points in the interval $-2\pi\mathfrak{i}[0,250]\cup 2\pi\mathfrak{i}[0,250]$ and symmetrically distributed according to eq. 30. We use the aforementioned methods to compute reduced-order models of dimension $r=75$. The resulting frequency responses and pointwise relative errors are plotted in Figure 4. Similar to the previous example, the structure-preserving approaches $\mathsf{soQuadpvBT}$ and $\mathsf{soLoewner}$ outperform the unstructured approach $\mathsf{foQuadBT}$ for most frequencies by about five orders of magnitude. The relative error measures reported in Table 4 also support this claim. We observe that all three reduced models struggle to produce a good approximation in the region around 48 Hz, which is due to the TVAs that heavily damp the frequency response about that region.
6.4 Mass-spring-damper network with velocity outputs
As the final example, we consider the mass-spring damper ($MSD$) chain from [43, Ex. 2]. This is a damped linear vibration system consisting of three rows of $d$ masses all connected to the right-hand side of a coupling mass $m_{0}$. The masses in the individual rows are connected by springs, and the coupling mass $m_{0}$ is connected to the fixed left-hand side of the base. The system order is $n=3d+1$. We construct a system with $d=300$ masses so that the full-order model dimension is $n=901$. We impose Rayleigh damping on the system for the parameters $f(s)=g(s)=0.002$, the input $\boldsymbol{B}_{\operatorname{u}}=\boldsymbol{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones, and we have the outputs $\boldsymbol{C}_{\operatorname{p}}=\boldsymbol{0}$ and $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{1}_{n}\in\mathbb{R}^{n}$. We include this example to illustrate how our method performs when velocity outputs are incorporated.
Figure 5: Frequency response and pointwise relative errors for $r=20$ reduced-order models of the coupled $MSD$ network with a velocity output. The proposed $\mathsf{soQuadpvBT}$ approach yields similar performance to its intrusive counterpart $\mathsf{soBT}$.
Table 5: Relative $\mathcal{H}_{\infty}$ errors eq. 36a and $\mathcal{H}_{2}$ errors eq. 36b for $r=20$ reduced-order models of the $MSD$ chain. The smallest error is highlighted in boldface.
$\mathsf{soQuadpvBT}$ $\mathsf{soLoewner}$ $\mathsf{foQuadBT}$ $\mathsf{sopvBT}$ $\mathsf{BT}$
[…]
We compute reduced-order models of size $r=20$ using $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$, $\mathsf{foQuadBT}$, $\mathsf{sopvBT}$ and $\mathsf{BT}$. For generating the data, we choose $N=200$ left and right quadrature nodes as interwoven logarithmically spaced points in the interval $-\mathfrak{i}[10^{-3},10^{1}]\cup\mathfrak{i}[10^{-3},10^{1}]$, which are symmetrically distributed according to eq. 30. The frequency response of the full and reduced-order transfer functions as well as their pointwise relative errors eq. 35 are plotted in Figure 5. All methods produce reasonably good approximations to the full-order system response. As was the case for the previous two benchmark problems, we again observe that the structure-preserving methods $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$, and $\mathsf{soBT}$ outperform the unstructured methods $\mathsf{foQuadBT}$ and $\mathsf{BT}$. This performance gap is particularly evident at the lower frequencies; all four methods exhibit roughly similar behavior from $1$ rad/s onward. The relative error measures are reported in Table 5, which shows that $\mathsf{sopvBT}$ performs the best overall.
7 Conclusions
We introduced a data-driven formulation of the second-order position-velocity balanced truncation approach for the construction of structured reduced-order systems solely from transfer function data. The approach is based on the quadrature formulation of the Gramian integrals used in the balancing process. We provide several modifications of the approach for practical situations including the preservation of stability in the case of symmetric second-order systems, the construction of real-valued systems, and the inference of optimal damping parameters from given data. The numerical experiments verify that the data-driven formulation of the position-velocity balanced truncation is indistinguishable from its intrusive counterpart for accurate enough quadrature rules, and they show the effectiveness of the approach to construct structured reduced-order models, which successfully outperform unstructured data-driven models of the same size.